### v1

# Carleman Estimates and Controllability of Forward Stochastic Parabolic Equations with General Dynamic Boundary Conditions

Abstract
We derive a new Carleman estimate for a general backward stochastic parabolic equation with dynamic boundary conditions, incorporating weak divergence source terms in both the bulk and surface equations. This estimate is obtained through two main steps: first, by refining a known Carleman estimate for backward stochastic parabolic equations to explicitly account for the dependence of the parameters on the final control time $T$; second, by applying a duality technique to address weak divergence source terms. As an application, we prove the null and approximate controllability of forward stochastic parabolic equations with dynamic boundary conditions, which involve both reaction and convection terms with bounded adapted coefficients, as well as general second-order parabolic operators. Additionally, we provide an explicit estimate for the null-controllability cost in terms of the final control time $T$ and the coefficients of the equation.
Cadi Ayyad University, National School of Applied Sciences, Laboratory of Mathematics, Modeling and Automatic Systems, B.P. 575, Marrakesh, Morocco. Email: s.boulite@uca.ma
Dipartimento di Matematica, Università degli Studi di Salerno, Via Giovanni Paolo II, 132, 84084 Fisciano (SA), Italy. Email: aelgrou@unisa.it,  arhandi@unisa.it
§Cadi Ayyad University, Faculty of Sciences Semlalia, Laboratory of Mathematics, Modeling and Automatic Systems, B.P. 2390, Marrakesh, Morocco. Email: maniar@uca.ma
University Mohammed VI Polytechnic, Vanguard Center, Benguerir, Morocco.
Email: Lahcen.Maniar@um6p.ma
Keywords: Carleman estimates; Stochastic parabolic equations; Null controllability; Approximate controllability; Observability inequality; Unique continuation property; Dynamic boundary conditions.
Mathematics Subject Classification: 93B05, 93B07, 93E20, 60H15.
1 Introduction and Main Results
Let $T>0$, $G\subset\mathbb{R}^{N}$ be a given nonempty open bounded domain with a smooth boundary $\Gamma=\partial G$, and $N\geq 2$. Let $G_{0}\Subset G$ (i.e., $\overline{G_{0}}\subset G$, where $\overline{G_{0}}$ denotes the closure of $G_{0}$) be a nonempty open subset strictly contained in $G$, with $\mathbbm{1}_{G_{0}}$ denoting the characteristic function of $G_{0}$. Define the following sets:
$Q=(0,T)\times G,\quad\Sigma=(0,T)\times\Gamma,\quad\text{and}\quad Q_{0}=(0,T)\times G_{0}.$
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a fixed complete filtered probability space on which a one-dimensional standard Brownian motion $W(\cdot)$ is defined, such that $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the natural filtration generated by $W(\cdot)$ and augmented by all the $\mathbb{P}$-null sets in $\mathcal{F}$. Let $H$ be a Banach space, and let $C([0,T];H)$ be the Banach space of all $H$-valued continuous functions defined on $[0,T]$. For some sub-sigma algebra $\mathcal{G}\subset\mathcal{F}$, we denote by $L^{2}_{\mathcal{G}}(\Omega;H)$ the Banach space of all $H$-valued $\mathcal{G}$-measurable random variables $X$ such that $\mathbb{E}\left[\|X\|_{H}^{2}\right]<\infty$, with the canonical norm. We also denote by $L^{2}_{\mathcal{F}}(0,T;H)$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted processes $X(\cdot)$ such that $\mathbb{E}\left[\|X(\cdot)\|^{2}_{L^{2}(0,T;H)}\right]<\infty$, with the canonical norm. Moreover, we denote by $L^{\infty}_{\mathcal{F}}(0,T;H)$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted essentially bounded processes. Finally, we denote by $L^{2}_{\mathcal{F}}(\Omega;C([0,T];H))$ the Banach space consisting of all $H$-valued $\{\mathcal{F}_{t}\}_{t\geq 0}$-adapted continuous processes $X(\cdot)$ such that $\mathbb{E}\left[\|X\|^{2}_{C([0,T];H)}\right]<\infty$, with the canonical norm.
Consider the following Hilbert space
$\mathbb{L}^{2}:=L^{2}(G,dx)\times L^{2}(\Gamma,d\sigma)=L^{2}(G)\times L^{2}(\Gamma),$
where $dx$ denotes the Lebesgue measure on $G$ and $d\sigma$ denotes the surface measure on $\Gamma$, equipped with the following inner product:
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{L}^{2}}=\langle y,z\rangle_{L^{2}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{L^{2}(\Gamma)}.$
We assume that the diffusion matrices $\mathcal{A}=(a^{jk})_{j,k=1,2,\cdots,N}$ and $\mathcal{A}_{\Gamma}=(b^{jk})_{j,k=1,2,\cdots,N}$ satisfy the following assumptions:
  1. 1.
    $\mathcal{A}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{2,\infty}(\overline{G};\mathbb{R}^{N\times N})))$, and $a^{jk}=a^{kj}$ for all $j,k=1,2,\cdots,N$.
  2. 2.
    $\mathcal{A}_{\Gamma}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{1,\infty}(\Gamma;\mathbb{R}^{N\times N})))$, and $b^{jk}=b^{kj}$ for all $j,k=1,2,\cdots,N$.
  3. 3.
    There exists a constant $\beta_{0}>0$ such that
    $\begin{aligned} \langle\mathcal{A}(\omega,t,x)\xi,\xi\rangle_{\mathbb{R}^{N}}&\geq\beta_{0}|\xi|^{2}\quad\text{for any}\quad(\omega,t,x,\xi)\in\Omega\times(0,T)\times\overline{G}\times\mathbb{R}^{N},\\ \langle\mathcal{A}_{\Gamma}(\omega,t,x)\xi,\xi\rangle_{\mathbb{R}^{N}}&\geq\beta_{0}|\xi|^{2}\quad\text{for any}\quad(\omega,t,x,\xi)\in\Omega\times(0,T)\times\Gamma\times\mathbb{R}^{N},\end{aligned}$ (1.1)
    where $\langle\cdot,\cdot\rangle_{\mathbb{R}^{N}}$ (resp. $|\cdot|$) is the Euclidean inner product (resp. norm) in $\mathbb{R}^{N}$.
Throughout this paper, $C$ denotes a positive constant depending only on $G$, $G_{0}$, $\mathcal{A}$, and $\mathcal{A}_{\Gamma}$, which may change from one place to another.
Consider the following non-homogeneous linear backward stochastic parabolic equation subject to dynamic boundary conditions
$\begin{cases}\begin{array}[]{ll}dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt=(F_{1}+\nabla\cdot F)\,dt+Z\,dW(t)&\text{in}\,Q,\\ dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt=(F_{2}-F\cdot\nu+\nabla_{\Gamma}\cdot F_{\Gamma})\,dt+\widehat{Z}\,dW(t)&\text{on}\,\Sigma,\\ z_{\Gamma}=z|_{\Gamma}&\text{on}\,\Sigma,\\ (z,z_{\Gamma})|_{t=T}=(z_{T},z_{\Gamma,T})&\text{in}\,G\times\Gamma,\end{array}\end{cases}$ (1.2)
[…]
$\partial^{\mathcal{A}}_{\nu}z=(\mathcal{A}\nabla z\cdot\nu)|_{\Sigma}=\sum_{j,k=1}^{N}a^{jk}\frac{\partial z}{\partial x_{j}}\nu^{k},$
represents the conormal derivative of $z$ with respect to $\mathcal{A}$, where $\nu=(\nu^{1},\cdots,\nu^{N})$ is the outer unit normal vector at $\Gamma$. This conormal derivative acts as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$, meaning the bulk equation is controlled directly, while the surface equation is governed through the conormal derivative $\partial_{\nu}^{\mathcal{A}}z$.
The operators $\nabla\cdot$ (in $G$) and $\nabla_{\Gamma}\cdot$ (on $\Gamma$) are the divergence operators. If $\mathcal{A}=I_{N}$, the identity matrix, then the conormal derivative $\partial_{\nu}^{\mathcal{A}}z$ simplifies to the standard normal derivative $\partial_{\nu}z$, and the bulk diffusion term $\nabla\cdot(\mathcal{A}\nabla z)$ becomes the Laplacian $\Delta z$. The surface term $\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})$ captures surface diffusion effects on $\Gamma$, where the tangential gradient $\nabla_{\Gamma}z_{\Gamma}:=\nabla z-(\partial_{\nu}z)\nu$ is used. In this case, if $\mathcal{A}_{\Gamma}=I_{N}$, the surface diffusion term simplifies to the Laplace-Beltrami operator $\Delta_{\Gamma}z_{\Gamma}$ on $\Gamma$. For further details on these geometric tools and on dynamic boundary conditions, we refer the reader to [17, 20, 28, 33].
Similarly to the standard Sobolev spaces $H^{k}(G)$ ($k=1,2$), we introduce the Sobolev spaces $H^{k}(\Gamma)$ ($k=1,2$) on the boundary $\Gamma$ as follows:
$H^{k}(\Gamma)=\left\{z_{\Gamma}\in L^{2}(\Gamma)\mid\nabla_{\Gamma}z_{\Gamma}\in H^{k-1}(\Gamma;\mathbb{R}^{N})\right\},\quad k=1,2,$
which are endowed with the following norms:
$|z_{\Gamma}|_{H^{1}(\Gamma)}=\sqrt{\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}},\quad\text{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{1}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\nabla_{\Gamma}y_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
and
$|z_{\Gamma}|_{H^{2}(\Gamma)}=\sqrt{\langle z_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}},\quad\text{where}\quad\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{2}(\Gamma)}=\int_{\Gamma}y_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}\Delta_{\Gamma}y_{\Gamma}\,\Delta_{\Gamma}z_{\Gamma}\,d\sigma.$
In the sequel, we also define the following Hilbert spaces:
$\mathbb{H}^{k}=\{(z,z_{\Gamma})\in H^{k}(G)\times H^{k}(\Gamma):z_{\Gamma}=z|_{\Gamma}\},\quad k=1,2,$
equipped with the usual inner product
$\langle(y,y_{\Gamma}),(z,z_{\Gamma})\rangle_{\mathbb{H}^{k}}=\langle y,z\rangle_{H^{k}(G)}+\langle y_{\Gamma},z_{\Gamma}\rangle_{H^{k}(\Gamma)}.$
In this paper, the following property
$(\textbf{z},\textbf{z}_{\Gamma})=(z_{1},z_{2})\quad\text{in}\;G\times\Gamma,\quad\textnormal{a.s.},$
is understood as follows:
$\textbf{z}=z_{1}\quad\text{in}\;G\quad\text{and}\quad\textbf{z}_{\Gamma}=z_{2}\quad\text{on}\;\Gamma,\quad\textnormal{a.s.}$
Remark 1.1.
In this paper, the concept of vectors on the boundary $\Gamma$ is understood in the context of vector fields on the manifold $\Gamma$. More precisely, these vectors are functions $\gamma:\Gamma\to\mathbb{R}^{N}$ such that for each $x\in\Gamma$, the vector $\gamma(x)$ lies in the tangent space $T_{x}\Gamma$ at $x$, defined by
$T_{x}\Gamma=\{p\in\mathbb{R}^{N}\mid p\cdot\nu(x)=0\}.$
This assumption allows for simplifications in the computation of certain boundary terms via the tangential divergence formula:
$\int_{\Gamma}(\nabla_{\Gamma}\cdot\gamma)\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}\gamma\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma,$
for $\gamma\in H^{1}(\Gamma;\mathbb{R}^{N})$ and $z_{\Gamma}\in H^{1}(\Gamma)$. However, when the vector field $\gamma:\Gamma\to\mathbb{R}^{N}$ is not necessarily tangential, one needs to use the generalized divergence formula (see [16, Proposition 5.4.9]):
$\int_{\Gamma}(\nabla_{\Gamma}\cdot\gamma)\,z_{\Gamma}\,d\sigma=-\int_{\Gamma}\gamma\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma+\int_{\Gamma}(\nabla_{\Gamma}\cdot\nu)\,(\gamma\cdot\nu)\,z_{\Gamma}\,d\sigma.$
This identity introduces additional boundary-like terms that need to be carefully treated.
By the duality method as in [4, Section 2], it can be shown that the equation (1.2) is well-posed. Specifically, for all given data $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, $F_{1}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))$, $F\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))$, $F_{2}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, and $F_{\Gamma}\in L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))$, the system (1.2) has a unique weak solution
$(z,z_{\Gamma};Z,\widehat{Z})\in\left(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\right)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}).$
Furthermore, there exists a constant $C=C(G,T,\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that
$\begin{aligned} &\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}+\|(Z,\widehat{Z})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})}\\ &\leq C\left(\|(z_{T},z_{\Gamma,T})\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})}+\|F_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|F\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G;\mathbb{R}^{N}))}\right.\\ &\hskip 28.45274pt\left.+\|F_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}+\|F_{\Gamma}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma;\mathbb{R}^{N}))}\right).\end{aligned}$
We recall the following well-known technical result due to Fursikov-Imanuvilov (see [15]).
Lemma 1.1.
For any nonempty open subset $G_{1}\Subset G$, there exists a function $\psi\in C^{4}(\overline{G})$ such that
$\psi>0\quad\textnormal{in}\;\,G;\qquad\psi=0\quad\textnormal{on}\;\,\Gamma;\qquad|\nabla\psi|>0\quad\textnormal{in}\;\,\overline{G\setminus G_{1}}.$
Since $|\nabla\psi|^{2}=|\nabla_{\Gamma}\psi|^{2}+|\partial_{\nu}\psi|^{2}$ on $\Gamma$, the function $\psi$ in Lemma 1.1 satisfies
$\nabla_{\Gamma}\psi=0,\qquad|\nabla\psi|=|\partial_{\nu}\psi|,\qquad\partial_{\nu}\psi\leq-c<0\quad\text{on}\;\,\Gamma,\quad\text{for some constant }\,c>0.$ (1.3)
For any parameters $\lambda,\mu\geq 1$, we choose the following weight functions:
$\alpha\equiv\alpha(t,x)=\frac{e^{\mu\psi(x)}-e^{2\mu|\psi|_{\infty}}}{t(T-t)},\quad\ell=\lambda\alpha,\quad\theta=e^{\ell},\quad\varphi\equiv\varphi(t,x)=\frac{e^{\mu\psi(x)}}{t(T-t)},$ (1.4)
for any $(t,x)\in(0,T)\times\overline{G}$. Moreover, the functions $\alpha$ and $\varphi$ are constant on the boundary $\Gamma$, so that
$\nabla_{\Gamma}\alpha=\nabla_{\Gamma}\varphi=0\quad\text{on}\;\,\Gamma.$
It is easy to verify that for some suitable constant $C=C(G)>0$, the following inequalities hold:
$\begin{aligned} &\,\varphi\geq CT^{-2},\qquad|\varphi_{t}|\leq CT\varphi^{2},\qquad|\varphi_{tt}|\leq CT^{2}\varphi^{3},\\ &|\alpha_{t}|\leq CTe^{2\mu|\psi|_{\infty}}\varphi^{2},\qquad|\alpha_{tt}|\leq CT^{2}e^{2\mu|\psi|_{\infty}}\varphi^{3},\end{aligned}$ (1.5)
for all $(t,x)\in(0,T)\times\overline{G}$. In the following, we choose the set $G_{1}$ from Lemma 1.1 such that $G_{1}\Subset G_{0}$.
The first main result of this paper is the following Carleman estimate for the solutions of (1.2).
Theorem 1.1.
[…]
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}\,dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}\,d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}\,dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}\,d\sigma dt\Bigg],\end{aligned}$ (1.6)
for all $\mu\geq\mu_{0}$ and $\lambda\geq\lambda_{0}(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
In [6, Theorem 4.1], a Carleman estimate for equation (1.2) is presented in the absence of divergence source terms (i.e., $F\equiv F_{\Gamma}\equiv 0$), where only Laplacian operators are involved (i.e., $\mathcal{A}\equiv\mathcal{A}_{\Gamma}\equiv I_{N}$), and where the parameters $\lambda$ and $\mu$ do not depend on the final control time $T$. This estimate is derived using the weighted identity method, applying two weighted identities for the stochastic parabolic operators $\text{``}dz+\Delta z\,dt\text{''}$ and $\text{``}dz_{\Gamma}+\Delta_{\Gamma}z_{\Gamma}\,dt-\partial_{\nu}z\,dt\text{''}$. In this paper, we extend this result to address the presence of divergence source terms and general second-order parabolic operators. Our goal is to derive the new Carleman estimate, given in Theorem 1.1. To obtain this estimate, we combine two key methods. First, we apply the weighted identity method to the operators $\text{``}dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt\text{''}$ and $\text{``}dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial^{\mathcal{A}}_{\nu}z\,dt\text{''}$, refining the known Carleman estimate from [6, Theorem 4.1] to a more general form. This generalization includes diffusion terms represented by the matrices $\mathcal{A}$ and $\mathcal{A}_{\Gamma}$, and explicitly incorporates the dependence of $\lambda$ and $\mu$ on the final control time $T$. Second, to handle the divergence source terms $F$ and $F_{\Gamma}$, we employ the duality technique introduced in [23] for the stochastic setting, which ultimately leads to the desired Carleman estimate (1.6). For more details on the derivation of Carleman estimates for stochastic parabolic equations with different types of boundary conditions, we refer the reader to [3, 14, 23, 32] and [27, Chapter 9] for the case of Dirichlet boundary conditions, to [7, 34] for the case of Neumann and Robin boundary conditions, and to [4, 5, 6] for the case of dynamic boundary conditions. For results on Carleman estimates for some coupled forward-backward stochastic parabolic equations, see [23, 29].
As an application of our Carleman estimate (1.6), we study the null and approximate controllability problems for the following forward stochastic parabolic equation
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u)\,dt+v_{1}\,dW(t)&\text{in}\,\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+v_{2}\,dW(t)&\text{on}\,\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\text{on}\,\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0})&\text{in}\,\,G\times\Gamma,\end{array}\end{cases}$ (1.7)
where $(y,y_{\Gamma})$ is the state variable, $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ is the initial state, the coefficients $a_{1}$, $a_{2}$, $B_{1}$, and $B_{2}$ are bounded and adapted stochastic processes, i.e.,
$a_{1}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G)),\quad B_{1}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(G;\mathbb{R}^{N})),$
$a_{2}$ $\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma)),\quad B_{2}\in L_{\mathcal{F}}^{\infty}(0,T;L^{\infty}(\Gamma;\mathbb{R}^{N})).$
In (1.7), the control function consists of the triple
$(u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)).$
From [4, Theorem 2.1], we deduce that (1.7) is well-posed, i.e., for any $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ and $(u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, the system (1.7) admits a unique weak solution
$(y,y_{\Gamma})\in L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1}),$
such that for some positive constant $C=C(G,T,\mathcal{A},\mathcal{A}_{\Gamma})$, we have
$\,\|(y,y_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(y,y_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}$
$\leq C\Big(\|(y_{0},y_{\Gamma,0})\|_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}+\|u\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))}$
$\hskip 28.45274pt+\|v_{1}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|v_{2}\|_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}\Big).$
The equation (1.7) with $u\equiv v_{1}\equiv v_{2}\equiv 0$ models diffusion phenomena, such as the evolution of quantities like heat, bacterial populations, or chemical concentrations, under stochastic perturbations and interactions with the boundary through the dynamic boundary conditions ((1.7)2 and (1.7)3), which involve the time derivative of the state on the boundary. Additionally, the convection terms $B_{1}\cdot\nabla y$ and $B_{2}\cdot\nabla_{\Gamma}y_{\Gamma}$ in (1.7) enable the modeling of these diffusion processes in a moving medium, such as fluid flow with a specific velocity. For further details on the model of stochastic diffusion equations, we refer to [10, 27] and the references therein.
The second main result of this paper is the following approximate controllability of (1.7).
Theorem 1.2.
Equation (1.7) is approximately controllable at time $T$ and for any control region $G_{0}\Subset G$. That is, for any initial state $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$ and final state $(y_{T},y_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, and for all $\varepsilon>0$, there exist controls
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ to (1.7) satisfies
$\mathbb{E}\|(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))-(y_{T},y_{\Gamma,T})\|^{2}_{\mathbb{L}^{2}}\leq\varepsilon.$
The third main result is the following null controllability of (1.7).
Theorem 1.3.
Equation (1.7) is null controllable at time $T$ and for any control region $G_{0}\Subset G$. That is, for any initial state $(y_{0},y_{\Gamma,0})\in L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})$, there exist controls
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)),$
such that the associated solution $(\widehat{y},\widehat{y}_{\Gamma})$ to (1.7) satisfies
$(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))=(0,0)\quad\textnormal{in}\;\,G\times\Gamma,\quad\textnormal{a.s.}$
Moreover, the controls $\widehat{u}$, $\widehat{v}_{1}$, and $\widehat{v}_{2}$ can be chosen such that
$\begin{aligned} \|\widehat{u}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))}+\|\widehat{v}_{1}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|\widehat{v}_{2}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}\leq e^{CK}\|(y_{0},y_{\Gamma,0})\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})},\end{aligned}$ (1.8)
where $K$ has the following form:
$K\equiv 1+\frac{1}{T}+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+T(\|a_{1}\|_{\infty}+\|a_{2}\|_{\infty})+(1+T)(\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}).$
Remark 1.2.
In (1.7), by acting on the system with the controls $u$, $v_{1}$, and $v_{2}$, we aim to influence the final behavior of the evolution. The conormal derivative $\partial_{\nu}^{\mathcal{A}}y$ serves as the coupling term between the bulk equation in $Q$ and the surface equation on $\Sigma$. Thus, the controls $u$, $v_{1}$, and $v_{2}$ directly affect the drift term through the control $u$ on the small region $G_{0}$ of the bulk equation, and the diffusion terms of both the bulk and surface equations via the extra controls $v_{1}$ and $v_{2}$, which act throughout the entire spatial domain $G$. Through this coupling term $\partial_{\nu}^{\mathcal{A}}y$, these controls also influence the drift term of the surface equation. This interplay allows us to establish the null and approximate controllability of (1.7).
This paper investigates the null and approximate controllability of a class of stochastic parabolic equations with dynamic boundary conditions. For results on the controllability of deterministic parabolic equations with such boundary conditions, we refer the reader to [28], [1], and [20]. The case of fully discrete parabolic equations is addressed in [21]. Additionally, numerous results on the controllability of stochastic partial differential equations are available in the literature; for a comprehensive discussion, we refer the reader to the book [27] and the references therein.
The controllability of stochastic parabolic equations with various boundary conditions—both static and dynamic—has been extensively studied. Specifically, for static boundary conditions, controllability has been investigated for Dirichlet, Neumann, and Robin boundary conditions. For Dirichlet boundary conditions, [2] was among the first to establish controllability results for forward stochastic parabolic equations with a localized control in the drift term, under certain restrictive assumptions on the system’s coefficients. Later, [32] extended these results using global Carleman estimates to address the controllability of general forward stochastic parabolic equations, introducing an additional control in the diffusion term and considering only bounded adapted zero-order coefficients that may depend on space, time, and random variables. We also refer to [14] where a general weighted identity for some stochastic operators with complex principal parts is established and then used to study controllability of certain type of stochastic partial differential equations as stochastic complex Ginzburg–Landau equations. In [32], the authors achieved these results by introducing a second control acting on the diffusion term across the entire domain. The work [3] further investigates the controllability of more general stochastic parabolic equations, incorporating both zero- and first-order terms, under the assumption that all coefficients are only bounded and adapted. In [24, 35], the Lebeau-Robiano strategy was applied to explore both approximate and null controllability of the forward stochastic heat equation with a single control in the drift term, assuming space-independent coefficients. Additionally, the controllability of some classes of stochastic parabolic equations with Robin boundary conditions has been studied in [7, 34]. For results on the null controllability of a class of stochastic semidiscrete parabolic equations, see [37].
For controllability results of stochastic parabolic equations with dynamic boundary conditions, we refer to [6] for forward and backward stochastic parabolic equations with only reaction terms, [5] for a class of one-dimensional stochastic heat equations, [4] for backward stochastic parabolic equations that incorporate both reaction and convection terms, and [8] for the case of the forward stochastic heat equation with a time-space localized control in the drift term of the bulk equation, without extra forces. For controllability results concerning certain coupled forward-backward stochastic parabolic equations, we refer the reader to [23, 29].
Following the classical duality argument, the controllability of linear stochastic partial differential equations, as explored in the previously cited works, can be proven through an observability problem for the corresponding adjoint equation. This involves unique continuation for approximate controllability and an observability inequality for null controllability. A key tool in proving these observability results (for both parabolic PDEs and SPDEs) is Carleman estimates, which provide energy estimates with exponential weights and parameters sufficiently large to absorb undesired lower-order terms. Carleman estimates, first introduced by Carleman in 1939 to prove the uniqueness of solutions for second-order elliptic partial differential equations with two variables ([9]), are essential not only for analyzing control problems (see [27]) but also for solving certain inverse problems (e.g., [25, 26]).
To the best of our knowledge, this paper is the first to investigate the null and approximate controllability of a more general class of linear forward stochastic parabolic equations with dynamic boundary conditions. This class includes both reaction and convection terms, as well as general second-order parabolic operators. Furthermore, we establish the estimate (1.8) for the null-controllability cost of (1.7), which is expressed in terms of the final control time $T$ and the coefficients $a_{1}$, $a_{2}$, $B_{1}$, and $B_{2}$.
Now, some remarks are in order.
Remark 1.3.
In equation (1.7), since the controls $v_{1}$ and $v_{2}$ influence the diffusion terms in the whole domain $G$ and on its boundary $\Gamma$, respectively, the controllability results derived for (1.7) can be generalized to cases involving the state $(y,y_{\Gamma})$ and the gradients $\nabla y$ and $\nabla_{\Gamma}y_{\Gamma}$ in the diffusion terms. Specifically, the diffusion terms $v_{1}\,dW(t)$ (resp., $v_{2}\,dW(t)$) can be replaced by the more general expressions $b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1}]\,dW(t)$ (resp., $b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2}]\,dW(t)$), where the coefficients $b_{i}$ and $\widetilde{B}_{i}$, for $i=1,2$, are adapted and bounded stochastic processes. Indeed, in the case where (1.7) is null (resp., approximately) controllable, one can find the control triple $(u,v_{1},v_{2})$ such that
$(y(T,\cdot),y_{\Gamma}(T,\cdot))=(0,0)\quad\textnormal{in}\;G\times\Gamma,\quad\textnormal{a.s.}$
It then follows that the new control triple $(u^{\star},v^{\star}_{1},v^{\star}_{2})$, defined by
$u^{\star}=u,\qquad v^{\star}_{1}=v_{1}-(b_{1}y+\widetilde{B}_{1}\cdot\nabla y),\qquad v^{\star}_{2}=v_{2}-(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}),$
achieves the null (resp., approximate) controllability for the following extended system:
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(a_{1}y+B_{1}\cdot\nabla y+\mathbbm{1}_{G_{0}}u^{\star})\,dt+(b_{1}y+\widetilde{B}_{1}\cdot\nabla y+v_{1}^{\star})\,dW(t)&\textnormal{in}\,\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+(b_{2}y_{\Gamma}+\widetilde{B}_{2}\cdot\nabla_{\Gamma}y_{\Gamma}+v_{2}^{\star})\,dW(t)&\textnormal{on}\,\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\textnormal{on}\,\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(y_{0},y_{\Gamma,0})&\textnormal{in}\,\,G\times\Gamma.\end{array}\end{cases}$
Remark 1.4.
The surface diffusion term $\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})$ on the boundary in (1.7) plays a crucial role in deriving the Carleman estimate (1.6), specifically in the proof of Lemma 2.1—Step 2 in Appendix A. This term is essential for absorbing the following problematic integral from the right-hand side:
$\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt.$
However, the study of the controllability of (1.7) with the following dynamic boundary condition:
$\begin{cases}dy_{\Gamma}+\partial_{\nu}^{\mathcal{A}}y\,dt=(a_{2}y_{\Gamma}+B_{2}\cdot\nabla_{\Gamma}y_{\Gamma})\,dt+v_{2}\,dW(t)&\textnormal{on}\;\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\textnormal{on}\;\Sigma,\end{cases}$
remains an open problem. In contrast, it is well known that for $N=1$, the controllability of stochastic parabolic equations can be studied under the above type of boundary condition; see [5].
Remark 1.5.
In [31], controllability results are provided for semilinear stochastic parabolic equations with Dirichlet boundary conditions and globally Lipschitz semilinear terms. Extending these results to equations with dynamic boundary conditions, where the semilinear terms involve gradient terms in both the bulk and the boundary, is a natural next step. However, for a broad class of nonlinearities, as discussed in [12, 13], this remains an open problem due to the lack of compactness embedding in stochastic partial differential equations.
Remark 1.6.
[…]
$\|z(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(G))}+\|z_{\Gamma}(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\Gamma))}\leq C\,\mathbb{E}\iint_{Q_{0}}z^{2}\,dxdt.$
Alternatively, one may establish the unique continuation property:
$z=0\quad\textnormal{in}\;Q_{0},\,\;\textnormal{a.s.}\quad\Longrightarrow\quad(z_{T},z_{\Gamma,T})=(0,0)\quad\textnormal{in}\;G\times\Gamma,\,\;\textnormal{a.s.}$
For partial results related to the above problem, we refer the reader to [8, 24]. Introducing the additional controls $v_{1}$ and $v_{2}$ in the diffusion terms of equation (1.7), as discussed in Section 3, allows us to prove only a weaker observability inequality (the inequality (3.8)) and a weak unique continuation property (the property (3.7)) for the adjoint system (3.1).
In this paper, we use the following lemma to perform integration by parts in space (see, e.g., [11, 19]).
Lemma 1.2.
For any $F\in L^{2}(G;\mathbb{R}^{N})$ and $F_{\Gamma}\in L^{2}(\Gamma;\mathbb{R}^{N})$, we define the extension of the operator divergence of $F$ (resp., tangential divergence of $F_{\Gamma}$) as the continuous linear operator on $H^{1}(G)$ (resp., on $H^{1}(\Gamma)$) as follows:
$\nabla\cdot F:H^{1}(G)\longrightarrow\mathbb{R},\quad\textbf{v}\longmapsto-\int_{G}F\cdot\nabla\textbf{v}\,dx+\langle F\cdot\nu,\textbf{v}|_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)},$
$\nabla_{\Gamma}\cdot F_{\Gamma}:H^{1}(\Gamma)\longrightarrow\mathbb{R},\quad\textbf{v}_{\Gamma}\longmapsto-\int_{\Gamma}F_{\Gamma}\cdot\nabla_{\Gamma}\textbf{v}_{\Gamma}\,d\sigma.$
Moreover, we have the following estimates:
$\left|\langle\nabla\cdot F,\textbf{v}\rangle_{(H^{1}(G))^{\prime},H^{1}(G)}\right|\leq C_{1}\|F\|_{L^{2}(G;\mathbb{R}^{N})}\|\textbf{v}\|_{H^{1}(G)},\quad\text{for all }\;F\in L^{2}(G;\mathbb{R}^{N}),\,\textbf{v}\in H^{1}(G),$
$\left|\langle\nabla_{\Gamma}\cdot F_{\Gamma},\textbf{v}_{\Gamma}\rangle_{H^{-1}(\Gamma),H^{1}(\Gamma)}\right|\leq C_{2}\|F_{\Gamma}\|_{L^{2}(\Gamma;\mathbb{R}^{N})}\|\textbf{v}_{\Gamma}\|_{H^{1}(\Gamma)},\quad\text{for all }\;F_{\Gamma}\in L^{2}(\Gamma;\mathbb{R}^{N}),\,\textbf{v}_{\Gamma}\in H^{1}(\Gamma),$
for some positive constants $C_{1}$ and $C_{2}$.
This paper is organized as follows. In Section 2, we establish the main Carleman estimate presented in Theorem 1.1. In Section 3, we address the necessary observability problems, including the unique continuation property and the observability inequality for the adjoint equation associated with (1.7). In Section 4, we use the classical duality argument to deduce the proofs of our main controllability results i.e., Theorems 1.2 and 1.3. Finally, in Appendix A, we complete the proof of the known Carleman estimate from [6, Theorem 4.1] by extending the Laplacian operators to our general second-order diffusion operators and carefully analyzing the dependence of the parameters on the final control time $T$. This estimate plays a crucial role in Section 2.
2 Carleman Estimate for Backward Stochastic Parabolic Equations: Proof of Theorem 1.1
This section is devoted to proving the main Carleman estimate in Theorem 1.1. To begin, we first establish the following Carleman estimate for system (1.2) without divergence source terms (i.e., with $F\equiv F_{\Gamma}\equiv 0$).
Lemma 2.1.
[…]
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}\,dxdt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}\,dxdt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\qquad\quad+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}\,dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}\,d\sigma dt\Bigg],\end{aligned}$ (2.1)
for large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
Since the proof of the above estimate (2.1) involves several technical steps, we include all details in Appendix A.
Based on the result of Lemma 2.1, we now proceed to prove Theorem 1.1. First, we establish a controllability result for the following controlled forward stochastic parabolic equation with dynamic boundary conditions and source terms:
$\begin{cases}\begin{array}[]{ll}dy-\nabla\cdot(\mathcal{A}\nabla y)\,dt=(\lambda^{3}\mu^{4}\theta^{2}\varphi^{3}z+\mathbbm{1}_{G_{0}}u)\,dt+v_{1}\,dW(t)&\text{in}\,Q,\\ dy_{\Gamma}-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\Gamma})\,dt+\partial_{\nu}^{\mathcal{A}}y\,dt=\lambda^{3}\mu^{3}\theta^{2}\varphi^{3}z_{\Gamma}\,dt+v_{2}\,dW(t)&\text{on}\,\Sigma,\\ y_{\Gamma}=y|_{\Gamma}&\text{on}\,\Sigma,\\ (y,y_{\Gamma})|_{t=0}=(0,0)&\text{in}\,G\times\Gamma,\end{array}\end{cases}$ (2.2)
[…]
Proposition 2.1.
There exists a triple of controls $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$ such that the corresponding solution $(\widehat{y},\widehat{y}_{\Gamma})$ of (2.2) satisfies
$(\widehat{y}(T,\cdot),\widehat{y}_{\Gamma}(T,\cdot))=(0,0)\quad\textnormal{in}\,\;G\times\Gamma,\,\;\textnormal{a.s.}$
Moreover, there exists a constant $C>0$, depending only on $G$, $G_{0}$, $\mathcal{A}$, and $\mathcal{A}_{\Gamma}$, such that the following inequality holds:
$\begin{aligned} &\,\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}\widehat{u}^{2}\,dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}\widehat{v}_{1}^{2}\,dxdt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}\widehat{v}_{2}^{2}\,d\sigma dt+\mathbb{E}\iint_{Q}\theta^{-2}\widehat{y}^{2}\,dxdt+\mathbb{E}\iint_{\Sigma}\theta^{-2}\widehat{y}_{\Gamma}^{2}\,d\sigma dt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}|\nabla\widehat{y}|^{2}\,dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}|\nabla_{\Gamma}\widehat{y}_{\Gamma}|^{2}\,d\sigma dt\\ &\leq C\bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}\,dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt\bigg],\end{aligned}$ (2.3)
for large $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
Proof.
For clarity, we present the proof in three steps.
Step 1. Fix $\varepsilon>0$ and define the functions:
$\theta_{\varepsilon}=e^{\lambda\alpha_{\varepsilon}},\qquad\alpha_{\varepsilon}\equiv\alpha_{\varepsilon}(t,x)=\frac{e^{\mu\psi(x)}-e^{2\mu|\psi|_{\infty}}}{(t+\varepsilon)(T-t+\varepsilon)}.$ (2.4)
It is easy to see that $\theta_{\varepsilon}\geq\theta$. Consider the following optimal control problem
$\inf\big\{J_{\varepsilon}(u,v_{1},v_{2})\,|\,(u,v_{1},v_{2})\in\mathcal{U}\big\},$ (2.5)
where
$J_{\varepsilon}(u,v_{1},v_{2})=$ $\,\frac{1}{2}\mathbb{E}\iint_{Q_{0}}\lambda^{-3}\mu^{-4}\theta^{-2}\varphi^{-3}u^{2}dxdt+\frac{1}{2}\mathbb{E}\iint_{Q}\lambda^{-2}\mu^{-2}\theta^{-2}\varphi^{-2}v_{1}^{2}dxdt$
$+\frac{1}{2}\mathbb{E}\iint_{\Sigma}\lambda^{-2}\mu^{-2}\theta^{-2}\varphi^{-2}v_{2}^{2}d\sigma dt+\frac{1}{2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y^{2}dxdt+\frac{1}{2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\Gamma}^{2}d\sigma dt$
$+\frac{1}{2\varepsilon}\mathbb{E}\int_{G}|y(T)|^{2}dx+\frac{1}{2\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\Gamma}(T)|^{2}d\sigma,$
and
$\mathcal{U}=\Bigg\{($ $u,v_{1},v_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma)):$
$\mathbb{E}\iint_{Q_{0}}\theta^{-2}\varphi^{-3}u^{2}dxdt<\infty\,,\quad\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{1}^{2}dxdt<\infty\,,\quad\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{2}^{2}d\sigma dt<\infty\Bigg\}.$
It is easy to see that $J_{\varepsilon}$ is well-defined, continuous, strictly convex, and coercive. Therefore, the problem (2.5) admits a unique optimal solution $(u_{\varepsilon},v_{\varepsilon,1},v_{\varepsilon,2})\in\mathcal{U}$. Moreover, by classical arguments: Euler-Lagrange equation and optimality system (see, e.g., [18, 22]), the optimal solution $(u_{\varepsilon},v_{\varepsilon,1},v_{\varepsilon,2})$ can be characterized by:
$u_{\varepsilon}=-\mathbbm{1}_{G_{0}}\lambda^{3}\mu^{4}\theta^{2}\varphi^{3}r_{\varepsilon}\,\,,\quad\quad v_{\varepsilon,1}=-\lambda^{2}\mu^{2}\theta^{2}\varphi^{2}R_{\varepsilon,1}\,\,,\quad\quad v_{\varepsilon,2}=-\lambda^{2}\mu^{2}\theta^{2}\varphi^{2}R_{\varepsilon,2},$ (2.6)
where $(r_{\varepsilon},r_{\varepsilon,\Gamma};R_{\varepsilon,1},R_{\varepsilon,2})$ is the solution of the backward equation
$\begin{cases}\begin{array}[]{ll}dr_{\varepsilon}+\nabla\cdot(\mathcal{A}\nabla r_{\varepsilon})\,dt=-\theta^{-2}_{\varepsilon}y_{\varepsilon}\,dt+R_{\varepsilon,1}\,dW(t)&\textnormal{in}\,\,Q,\\ dr_{\varepsilon,\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}r_{\varepsilon,\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}r_{\varepsilon}dt=-\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}\,dt+R_{\varepsilon,2}\,dW(t)&\textnormal{on}\,\,\Sigma,\\ r_{\varepsilon,\Gamma}(t,x)=r_{\varepsilon}|_{\Gamma}(t,x)&\textnormal{on}\,\,\Sigma,\\ \big(r_{\varepsilon},r_{\varepsilon,\Gamma}\big)|_{{t=T}}=(\frac{1}{\varepsilon}y_{\varepsilon}(T),\frac{1}{\varepsilon}y_{\varepsilon,\Gamma}(T))&\textnormal{in}\,\,G\times\Gamma,\end{array}\end{cases}$ (2.7)
with $(y_{\varepsilon},y_{\varepsilon,\Gamma})$ is the solution of (2.2) associated to the controls $(u_{\varepsilon},v_{\varepsilon,1},v_{\varepsilon,2})$.
Step 2. Differentiating $d\langle(y_{\varepsilon},y_{\varepsilon,\Gamma}),(r_{\varepsilon},r_{\varepsilon,\Gamma})\rangle_{\mathbb{L}^{2}}$ by Itô’s formula, integrating the equality on $(0,T)$ and recalling (2.6), we obtain that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &=\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}zr_{\varepsilon}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}r_{\varepsilon,\Gamma}d\sigma dt.\end{aligned}$ (2.8)
By Young’s inequality, we get from (2.8) that for all $\rho>0$,
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq\rho\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}r_{\varepsilon,\Gamma}^{2}d\sigma\Bigg]\\ &\;\;\;+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.9)
Using the Carleman estimate (2.1) for solutions of (2.7) and noting that $\theta\theta_{\varepsilon}^{-1}\leq 1$, then there exists a constant $C=C(G,G_{0},\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that for all $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, the inequality (2.9) implies that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\rho\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\qquad\;\quad+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\Bigg]\\ &\;\;\;+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.10)
Choosing a small enough $\rho>0$ in (2.10), we obtain that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}r_{\varepsilon}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}R_{\varepsilon,1}^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}R_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$
Recalling (2.6), it follows that
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}u_{\varepsilon}^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.11)
On the other hand, by computing $d(\theta^{-2}_{\varepsilon}\varphi^{-2}y_{\varepsilon}^{2})$ with Itô’s formula, integrating the equality on $Q$ and taking the expectation on both sides, we find that
 $\begin{aligned} &\,2\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}\mathcal{A}\nabla y_{\varepsilon}\cdot\nabla y_{\varepsilon}dxdt\\ &=-2\mathbb{E}\iint_{Q}y_{\varepsilon}\mathcal{A}\nabla y_{\varepsilon}\cdot\nabla(\theta_{\varepsilon}^{-2}\varphi^{-2})dxdt+2\mathbb{E}\int_{0}^{T}\langle\partial_{\nu}^{\mathcal{A}}y_{\varepsilon},\theta_{\varepsilon}^{-2}\varphi^{-2}y_{\varepsilon,\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt\\ &\quad\,+\mathbb{E}\iint_{Q}(\theta^{-2}_{\varepsilon}\varphi^{-2})_{t}y_{\varepsilon}^{2}dxdt+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\theta^{2}\varphi y_{\varepsilon}zdxdt\\ &\quad\,+2\mathbb{E}\iint_{Q_{0}}\theta^{-2}_{\varepsilon}\varphi^{-2}y_{\varepsilon}u_{\varepsilon}dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt.\end{aligned}$ (2.12)
Similarly to (2.12), by computing $d(\theta^{-2}_{\varepsilon}\varphi^{-2}y_{\varepsilon,\Gamma}^{2})$, we get
 $\begin{aligned} &\,2\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}\mathcal{A}_{\Gamma}\nabla_{\Gamma}y_{\varepsilon,\Gamma}\cdot\nabla_{\Gamma}y_{\varepsilon,\Gamma}d\sigma dt\\ &=-2\mathbb{E}\int_{0}^{T}\langle\partial_{\nu}^{\mathcal{A}}y_{\varepsilon},\theta_{\varepsilon}^{-2}\varphi^{-2}y_{\varepsilon,\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt\\ &\quad\,+\mathbb{E}\iint_{\Sigma}(\theta^{-2}_{\varepsilon}\varphi^{-2})_{t}y_{\varepsilon,\Gamma}^{2}d\sigma dt+2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\theta^{2}\varphi y_{\varepsilon,\Gamma}z_{\Gamma}d\sigma dt\\ &\quad\,+\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma.\end{aligned}$ (2.13)
By using (1.5) and (2.4), it is easy to check that for a large enough $\lambda\geq CT^{2}$, we have for any $(t,x)\in(0,T)\times\overline{G}$,
$|(\theta^{-2}_{\varepsilon}\varphi^{-2})_{t}|\leq CT\lambda e^{2\mu|\psi|_{\infty}}\theta_{\varepsilon}^{-2},\qquad|\nabla(\theta^{-2}_{\varepsilon}\varphi^{-2})|\leq C\lambda\mu\theta_{\varepsilon}^{-2}\varphi^{-1}.$ (2.14)
Recalling (1.1) and combining (2.12), (2.13) and (2.14), we conclude that for a large $\lambda\geq CT^{2}$, we obtain
 $\begin{aligned} &\,2\beta_{0}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda\mu\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}\varphi^{-1}|y_{\varepsilon}||\nabla y_{\varepsilon}|dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt\\ &\quad\,+2\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\quad\,+2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.15)
Applying Young’s inequality on the right-hand side of (2.15), we deduce that for any $\rho>0$,
 $\begin{aligned} &\,2\beta_{0}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq\rho\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\frac{C}{\rho}\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt\\ &\quad\,+2\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt+2\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.16)
Absorbing the first term on the right-hand side of (2.16) with the first term on the left-hand side by choosing a small $\rho>0$. Thus, multiplying the obtained inequality by $\lambda^{-2}\mu^{-2}$, we find that
$\begin{aligned} &\,\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon}||z|dxdt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q_{0}}\theta^{-1}_{\varepsilon}\theta^{-1}\varphi^{-2}|y_{\varepsilon}||u_{\varepsilon}|dxdt+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &\quad\,+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{-1}_{\varepsilon}\theta\varphi|y_{\varepsilon,\Gamma}||z_{\Gamma}|d\sigma dt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.17)
Using Young’s inequality for (2.17), we have that
$\begin{aligned} &\,\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt\\ &\quad\,+C\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt+C\lambda^{-4}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-4}u_{\varepsilon}^{2}dxdt+C\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}y_{\varepsilon}^{2}dxdt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt+CT\lambda^{-1}\mu^{-2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$
Taking a large enough $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$ and $\mu\geq C$, it follows that
$\begin{aligned} &\,\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}y_{\varepsilon,\Gamma}^{2}d\sigma dt+C\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+C\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}u_{\varepsilon}^{2}dxdt\\ &\quad\,+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt+C\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}v_{\varepsilon,2}^{2}d\sigma dt.\end{aligned}$ (2.18)
Combining (2.18) and (2.11), we end up with
$\begin{aligned} &\,\frac{1}{\varepsilon}\mathbb{E}\int_{G}|y_{\varepsilon}(T)|^{2}dx+\frac{1}{\varepsilon}\mathbb{E}\int_{\Gamma}|y_{\varepsilon,\Gamma}(T)|^{2}d\sigma+\mathbb{E}\iint_{Q}\theta_{\varepsilon}^{-2}y_{\varepsilon}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta_{\varepsilon}^{-2}y_{\varepsilon,\Gamma}^{2}d\sigma dt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla y_{\varepsilon}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}_{\varepsilon}\varphi^{-2}|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2}d\sigma dt\\ &+\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}u_{\varepsilon}^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}v_{\varepsilon,1}^{2}dxdt\\ &+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-1}v_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$ (2.19)
Step 3. Using the uniform estimate (2.19) and recalling (1.4) and (2.4), we have that
$\begin{aligned} &\,\mathbb{E}\iint_{Q}(y_{\varepsilon}^{2}+|\nabla y_{\varepsilon}|^{2})dxdt+\mathbb{E}\iint_{\Sigma}(y_{\varepsilon,\Gamma}^{2}+|\nabla_{\Gamma}y_{\varepsilon,\Gamma}|^{2})d\sigma dt\\ &+\mathbb{E}\iint_{Q}(u_{\varepsilon}^{2}+v_{\varepsilon,1}^{2})dxdt+\mathbb{E}\iint_{\Sigma}v_{\varepsilon,2}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg],\end{aligned}$
which implies that
$\begin{aligned} &\,\mathbb{E}\int_{0}^{T}\|u_{\varepsilon}\|_{L^{2}(G)}^{2}dt+\mathbb{E}\int_{0}^{T}\|v_{\varepsilon,1}\|_{L^{2}(G)}^{2}dt+\mathbb{E}\int_{0}^{T}\|v_{\varepsilon,2}\|_{L^{2}(\Gamma)}^{2}dt\\ &+\mathbb{E}\int_{0}^{T}\|y_{\varepsilon}\|_{H^{1}(G)}^{2}dt+\mathbb{E}\int_{0}^{T}\|y_{\varepsilon,\Gamma}\|^{2}_{H^{1}(\Gamma)}dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\Bigg].\end{aligned}$
Then, we deduce that there exists
$(\widehat{u},\widehat{v}_{1},\widehat{v}_{2},\widehat{y},\widehat{y}_{\Gamma})\in$ $\,\,L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$
$\times L^{2}_{\mathcal{F}}(0,T;H^{1}(G))\times L^{2}_{\mathcal{F}}(0,T;H^{1}(\Gamma)),$
such that as $\varepsilon\rightarrow 0$,
$u_{\varepsilon}\longrightarrow\widehat{u},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;L^{2}(G_{0}));$
$v_{\varepsilon,1}\longrightarrow\widehat{v}_{1},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;L^{2}(G));$
$v_{\varepsilon,2}\longrightarrow\widehat{v}_{2},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;L^{2}(\Gamma));$ (2.20)
$y_{\varepsilon}\longrightarrow\widehat{y},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;H^{1}(G));$
$y_{\varepsilon,\Gamma}\longrightarrow\widehat{y}_{\Gamma},\,\,\,\textnormal{weakly in}\,\,\,L^{2}((0,T)\times\Omega;H^{1}(\Gamma)).$
Now, we claim that $(\widehat{y},\widehat{y}_{\Gamma})$ is the solution of (2.2) associated to the controls $\widehat{u},\widehat{v}_{1}$ and $\widehat{v}_{2}$. To prove this fact, we suppose $(\widetilde{y},\widetilde{y}_{\Gamma}))$ is the unique solution in $L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\cap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})$ to (2.2) with controls $\widehat{u}$, $\widehat{v}_{1}$ and $\widehat{v}_{2}$. For any processes $(f,g)\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$, we consider the following backward stochastic parabolic equation
$\begin{cases}\begin{array}[]{ll}d\phi+\nabla\cdot(\mathcal{A}\nabla\phi)\,dt=f\,dt+\Phi\,dW(t)&\textnormal{in}\,\,Q,\\ d\phi_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}\phi_{\Gamma})dt-\partial_{\nu}^{\mathcal{A}}\phi\,dt=g\,dt+\widehat{\Phi}\,dW(t)&\textnormal{on}\,\,\Sigma,\\ \phi_{\Gamma}(t,x)=\phi|_{\Gamma}(t,x)&\textnormal{on}\,\,\Sigma,\\ (\phi,\phi_{\Gamma})|_{t=T}=(0,0)&\textnormal{in}\,\,G\times\Gamma.\end{array}\end{cases}$
By Itô’s formula, computing $d\langle(\phi,\phi_{\Gamma}),(\widetilde{y},\widetilde{y}_{\Gamma})\rangle_{\mathbb{L}^{2}}-d\langle(\phi,\phi_{\Gamma}),(y_{\varepsilon},y_{\varepsilon,\Gamma})\rangle_{\mathbb{L}^{2}}$”, integrating the obtained equality on $(0,T)$, then we deduce that
 $\begin{aligned} &\,0=\mathbb{E}\iint_{Q}f(y_{\varepsilon}-\widetilde{y})dxdt+\mathbb{E}\iint_{Q_{0}}\phi(u_{\varepsilon}-\widehat{u})dxdt+\mathbb{E}\iint_{Q}\Phi(v_{\varepsilon,1}-\widehat{v}_{1})dxdt\\ &\quad\;+\mathbb{E}\iint_{\Sigma}g(y_{\varepsilon,\Gamma}-\widetilde{y}_{\Gamma})d\sigma dt+\mathbb{E}\iint_{\Sigma}\widehat{\Phi}(v_{\varepsilon,2}-\widehat{v}_{2})d\sigma dt.\end{aligned}$ (2.21)
Letting $\varepsilon\rightarrow 0$ in (2.21), we get
$\begin{aligned} \mathbb{E}\iint_{Q}f(\widehat{y}-\widetilde{y})dxdt+\mathbb{E}\iint_{\Sigma}g(\widehat{y}_{\Gamma}-\widetilde{y}_{\Gamma})d\sigma dt=0.\end{aligned}$ (2.22)
From (2.22), we conclude that $\widetilde{y}=\widehat{y}$ in $Q$, a.s., and $\widetilde{y}_{\Gamma}=\widehat{y}_{\Gamma}$ on $\Sigma$, a.s. Then we deduce that $(\widehat{y},\widehat{y}_{\Gamma})$ is the solution of (2.2) with controls $\widehat{u},\widehat{v}_{1}$ and $\widehat{v}_{2}$. Finally, combining the uniform estimate (2.19) and the convergence results (2.20), we obtain the null controllability of (2.2), along with the desired inequality (2.3). This concludes the proof of Proposition 2.1. ∎
Now, we are in a position to prove Theorem 1.1.
Proof of Theorem 1.1.
[…]
$\begin{aligned} \lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt=&-\mathbb{E}\iint_{Q}[\mathbbm{1}_{G_{0}}\widehat{u}z+F_{1}\widehat{y}-F\cdot\nabla\widehat{y}+Z\widehat{v}_{1}]dxdt\\ &-\mathbb{E}\iint_{\Sigma}[F_{2}\widehat{y}_{\Gamma}-F_{\Gamma}\cdot\nabla_{\Gamma}\widehat{y}_{\Gamma}+\widehat{Z}\widehat{v}_{2}]d\sigma dt.\end{aligned}$ (2.23)
Applying Young’s inequality on the right-hand side of (2.23), we get that for all $\rho>0$,
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\leq\rho\Bigg[\lambda^{-3}\mu^{-4}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-3}\widehat{u}^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}\widehat{v}_{1}^{2}dxdt\\ &\hskip 22.76228pt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}\widehat{v}_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{-2}\widehat{y}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{-2}\widehat{y}_{\Gamma}^{2}d\sigma dt\\ &\hskip 22.76228pt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{-2}\varphi^{-2}|\nabla\widehat{y}|^{2}dxdt+\lambda^{-2}\mu^{-2}\mathbb{E}\iint_{\Sigma}\theta^{-2}\varphi^{-2}|\nabla_{\Gamma}\widehat{y}_{\Gamma}|^{2}d\sigma dt\Bigg]\\ &\quad+\frac{1}{4\rho}\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt\\ &\hskip 45.52458pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\hskip 45.52458pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt\Bigg].\end{aligned}$ (2.24)
Using Carleman estimate (2.3) and taking a small enough $\rho>0$, (2.24) implies that
$\begin{aligned} &\,\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\hskip 28.45274pt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt\Bigg],\end{aligned}$ (2.25)
for all $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$.
On the other hand, computing $d|\theta\varphi^{\frac{1}{2}}z|^{2}_{L^{2}(G)}$ by using Itô’s formula, integrating the equality with respect to $t\in(0,T)$ and taking the expectation on both sides, we obtain that
$\begin{aligned} 0=&\,\mathbb{E}\iint_{Q}(\theta^{2}\varphi)_{t}z^{2}dxdt+2\mathbb{E}\iint_{Q}\mathcal{A}\nabla z\cdot\nabla(\theta^{2}\varphi z)dxdt\\ &-2\mathbb{E}\int_{0}^{T}\langle\partial_{\nu}^{\mathcal{A}}z,\theta^{2}\varphi z_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt+2\mathbb{E}\iint_{Q}\theta^{2}\varphi zF_{1}\,dxdt\\ &-2\mathbb{E}\iint_{Q}F\cdot\nabla(\theta^{2}\varphi z)dxdt+2\mathbb{E}\int_{0}^{T}\langle F\cdot\nu,\theta^{2}\varphi z_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt\\ &+\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt.\end{aligned}$ (2.26)
Similarly to (2.26), by computing $d|\theta\varphi^{\frac{1}{2}}z_{\Gamma}|^{2}_{L^{2}(\Gamma)}$, integrating the result on $(0,T)$ and taking the expectation, we get that
$\begin{aligned} 0=&\,\mathbb{E}\iint_{\Sigma}(\theta^{2}\varphi)_{t}z_{\Gamma}^{2}d\sigma dt+2\mathbb{E}\iint_{\Sigma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}(\theta^{2}\varphi z_{\Gamma})d\sigma dt\\ &+2\mathbb{E}\int_{0}^{T}\langle\partial_{\nu}^{\mathcal{A}}z,\theta^{2}\varphi z_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}F_{2}\,d\sigma dt\\ &-2\mathbb{E}\int_{0}^{T}\langle F\cdot\nu,\theta^{2}\varphi z_{\Gamma}\rangle_{H^{-1/2}(\Gamma),H^{1/2}(\Gamma)}dt-2\mathbb{E}\iint_{\Sigma}F_{\Gamma}\cdot\nabla_{\Gamma}(\theta^{2}\varphi z_{\Gamma})d\sigma dt\\ &+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.27)
Combining (2.26) and (2.27), it follows that
$\begin{aligned} 0=&\,2\mathbb{E}\iint_{Q}\mathcal{A}\nabla z\cdot\nabla(\theta^{2}\varphi z)dxdt+2\mathbb{E}\iint_{\Sigma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}(\theta^{2}\varphi z_{\Gamma})d\sigma dt\\ &+\mathbb{E}\iint_{Q}(\theta^{2}\varphi)_{t}z^{2}dxdt+\mathbb{E}\iint_{\Sigma}(\theta^{2}\varphi)_{t}z_{\Gamma}^{2}d\sigma dt+2\mathbb{E}\iint_{Q}\theta^{2}\varphi zF_{1}\,dxdt\\ &+2\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}F_{2}\,d\sigma dt-2\mathbb{E}\iint_{Q}F\cdot\nabla(\theta^{2}\varphi z)dxdt\\ &-2\mathbb{E}\iint_{\Sigma}F_{\Gamma}\cdot\nabla_{\Gamma}(\theta^{2}\varphi z_{\Gamma})d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt,\end{aligned}$ (2.28)
which implies that
 $\begin{aligned} &\,2\mathbb{E}\iint_{Q}\theta^{2}\varphi\mathcal{A}\nabla z\cdot\nabla zdxdt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt\\ &=-2\mathbb{E}\iint_{Q}z\mathcal{A}\nabla z\cdot\nabla(\theta^{2}\varphi)dxdt-\mathbb{E}\iint_{Q}(\theta^{2}\varphi)_{t}z^{2}dxdt-\mathbb{E}\iint_{\Sigma}(\theta^{2}\varphi)_{t}z_{\Gamma}^{2}d\sigma dt\\ &\quad\,-2\mathbb{E}\iint_{Q}\theta^{2}\varphi zF_{1}\,dxdt-2\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}F_{2}\,d\sigma dt+2\mathbb{E}\iint_{Q}\theta^{2}\varphi F\cdot\nabla zdxdt\\ &\quad\,+2\mathbb{E}\iint_{Q}zF\cdot\nabla(\theta^{2}\varphi)dxdt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi F_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt\\ &\quad\,-\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt-\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.29)
It is easy to see that for large $\lambda\geq CT^{2}$, we have
$|(\theta^{2}\varphi)_{t}|\leq CT\lambda e^{2\mu|\psi|_{\infty}}\theta^{2}\varphi^{3}\quad\;\;\textnormal{and}\quad\;\;|\nabla(\theta^{2}\varphi)|\leq C\lambda\mu\theta^{2}\varphi^{2}.$ (2.30)
Recalling (1.1) and using (2.30), we obtain for large $\lambda\geq CT^{2}$,
$\begin{aligned} &\,\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda\mu\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|z||\nabla z|dxdt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi|z||F_{1}|\,dxdt\\ &\quad\,+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||F_{2}|\,d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi|F||\nabla z|dxdt\\ &\quad\,+C\lambda\mu\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|z||F|dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt\\ &\quad\,+C\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.31)
Applying Young’s inequality in the right-hand side of (2.31), we get for any $\varepsilon>0$,
$\begin{aligned} &\,\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq\varepsilon\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\varepsilon\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt+\frac{C}{\varepsilon}\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+\frac{C}{\varepsilon}\mathbb{E}\iint_{Q}\theta^{2}\varphi|F|^{2}dxdt+\frac{C}{\varepsilon}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}|^{2}d\sigma dt+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt\\ &\quad\,+CT\lambda e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda^{-1}\mu^{-2}\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}\,dxdt\\ &\quad\,+C\lambda^{-1}\mu^{-1}\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt+C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}\,d\sigma dt+C\lambda\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt\\ &\quad\,+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt+C\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.32)
Choosing a small enough $\varepsilon$ in (2.32), and multiplying the obtained inequality by $\lambda\mu^{2}$, we deduce that
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|F|^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|F_{\Gamma}|^{2}d\sigma dt\\ &\quad\,+CT\lambda^{2}\mu^{2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+CT\lambda^{2}\mu^{2}e^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}\,dxdt+C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}\,d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt+C\lambda^{2}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}z^{2}dxdt\\ &\quad\,+C\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi Z^{2}dxdt+C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.33)
Choosing a sufficiently large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$ in (2.33), we conclude that
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\lambda^{3}\mu^{4}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+C\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\,d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|F|^{2}dxdt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|F_{\Gamma}|^{2}d\sigma dt+C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}\,dxdt+C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}\,d\sigma dt\\ &\quad\,+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (2.34)
Finally, by combining (2.34) and (2.25), we obtain the desired Carleman estimate (1.6). This completes the proof of Theorem 1.1. ∎
3 Observability Problems
This section is addressed to proving the appropriate unique continuation property and observability inequality for the following adjoint backward equation of (1.7) in the sense introduced in [27, Chap. 7, Sect. 7.3]:
$\begin{cases}\begin{array}[]{ll}dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt=(-a_{1}z+\nabla\cdot(zB_{1}))\,dt+Z\,dW(t)&\textnormal{in}\,\,Q,\\ dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt=(-a_{2}z_{\Gamma}-zB_{1}\cdot\nu+\nabla_{\Gamma}\cdot(z_{\Gamma}B_{2}))\,dt+\widehat{Z}\,dW(t)&\textnormal{on}\,\,\Sigma,\\ z_{\Gamma}=z|_{\Gamma}&\textnormal{on}\,\,\Sigma,\\ (z,z_{\Gamma})|_{t=T}=(z_{T},z_{\Gamma,T})&\textnormal{in}\,\,G\times\Gamma,\end{array}\end{cases}$ (3.1)
where $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$ is the terminal state. Similar to the well-posedness of (1.2), we also have that (3.1) is well-posed i.e., for all $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, the system (3.1) admits a unique weak solution
$(z,z_{\Gamma};Z,\widehat{Z})\in\Big(L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))\bigcap L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})\Big)\times L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2}),$
such that
$\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(\Omega;C([0,T];\mathbb{L}^{2}))}+\|(z,z_{\Gamma})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{H}^{1})}+\|(Z,\widehat{Z})\|_{L^{2}_{\mathcal{F}}(0,T;\mathbb{L}^{2})}\leq C\|(z_{T},z_{\Gamma,T})\|_{L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})}.$ (3.2)
From Theorem 1.1, we deduce the following Carleman estimate for solutions of the system (3.1).
Corollary 3.1.
[…]
$\begin{aligned} &\,\lambda^{3}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &+\lambda\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}Z^{2}dxdt+\lambda^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\widehat{Z}\,^{2}d\sigma dt\Bigg],\end{aligned}$ (3.3)
for sufficiently large
$\lambda\geq\lambda_{0}\left[T+T^{2}(1+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2})\right].$
Proof.
By applying Carleman estimate (1.6) with
$F_{1}=-a_{1}z,\quad F=zB_{1},\quad F_{2}=-a_{2}z_{\Gamma},\quad F_{\Gamma}=z_{\Gamma}B_{2},$
we have that for $\lambda\geq C(T+T^{2})$,
$\begin{aligned} &\,\lambda^{3}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\lambda^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &+\lambda\mathbb{E}\iint_{Q}\theta^{2}\varphi|\nabla z|^{2}dxdt+\lambda\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\Bigg[\lambda^{3}\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\mathbb{E}\iint_{Q}\theta^{2}|a_{1}z|^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}|a_{2}z_{\Gamma}|^{2}d\sigma dt\\ &\qquad\;+\lambda^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|zB_{1}|^{2}dxdt+\lambda^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}B_{2}|^{2}d\sigma dt\\ &\qquad\;+\lambda^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+\lambda^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}\,^{2}d\sigma dt\Bigg].\end{aligned}$ (3.4)
First, it is easy to see that
$\begin{aligned} &C\mathbb{E}\iint_{Q}\theta^{2}|a_{1}z|^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|a_{2}z_{\Gamma}|^{2}d\sigma dt\\ &+C\lambda^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}|zB_{1}|^{2}dxdt+C\lambda^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}B_{2}|^{2}d\sigma dt\\ &\leq\frac{1}{2}\lambda^{3}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}z^{2}dxdt+\frac{1}{2}\lambda^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt,\end{aligned}$ (3.5)
for large $\lambda\geq CT^{2}(\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2})$.
We also have that for a large $\lambda\geq CT^{2}$,
$\begin{aligned} &C\lambda^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+C\lambda^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}\,^{2}d\sigma dt\\ &\leq\lambda^{3}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}Z^{2}dxdt+\lambda^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\widehat{Z}\,^{2}d\sigma dt.\end{aligned}$ (3.6)
Combining (3.4), (3.5) and (3.6), we deduce the desired estimate (3.3). ∎
From Corollary 3.1, we first deduce the following unique continuation property for the system (3.1).
Proposition 3.1.
[…]
$(z,Z,\widehat{Z})=(0,0,0)\,\,\,\textnormal{in}\,\,\,Q_{0}\times Q\times\Sigma,\;\,\textnormal{a.s.}\Longrightarrow\;\,(z_{T},z_{\Gamma,T})=(0,0),\,\,\,\textnormal{in}\,\,\,G\times\Gamma,\;\,\textnormal{a.s.}$ (3.7)
Secondly, from Corollary 3.1, we also derive the following observability inequality.
Proposition 3.2.
[…]
$\begin{aligned} &\,\|z(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(G))}+\|z_{\Gamma}(0,\cdot)\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;L^{2}(\Gamma))}\\ &\leq e^{CK}\,\bigg[\mathbb{E}\iint_{Q_{0}}z^{2}dxdt+\mathbb{E}\iint_{Q}Z^{2}dxdt+\mathbb{E}\iint_{\Sigma}\widehat{Z}\,^{2}d\sigma dt\bigg],\end{aligned}$ (3.8)
where $K$ has the following form:
$K\equiv 1+\frac{1}{T}+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+T(\|a_{1}\|_{\infty}+\|a_{2}\|_{\infty})+(1+T)(\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}).$
Proof.
From the Carleman inequality (3.3), we deduce
$\begin{aligned} &\,\mathbb{E}\int_{T/4}^{3T/4}\int_{G}\theta^{2}\varphi^{3}z^{2}dxdt+\mathbb{E}\int_{T/4}^{3T/4}\int_{\Gamma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\leq C\bigg[\mathbb{E}\iint_{Q_{0}}\theta^{2}\varphi^{3}z^{2}dxdt+\mathbb{E}\iint_{Q}\theta^{2}\varphi^{3}Z^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\widehat{Z}\,^{2}d\sigma dt\bigg],\end{aligned}$ (3.9)
for large $\lambda\geq\lambda_{1}(T)\equiv\lambda_{0}\left[T+T^{2}(1+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2})\right].$ Subsequently, notice that the functions
$t\mapsto\min_{x\in\overline{G}}\,[\theta^{2}(t,x)\varphi^{3}(t,x)]$
reach their minimum in $[T/4,3T/4]$ at $t=T/4$ and
$t\mapsto\max_{x\in\overline{G}}\,[\theta^{2}(t,x)\varphi^{3}(t,x)]$
reach their maximum in $(0,T)$ at $t=T/2$. Then, fixing $\lambda=\lambda_{1}(T)$, we deduce that
$\begin{aligned} &\,\mathbb{E}\int_{T/4}^{3T/4}\int_{G}z^{2}dxdt+\mathbb{E}\int_{T/4}^{3T/4}\int_{\Gamma}z_{\Gamma}^{2}d\sigma dt\\ &\leq e^{C\textbf{K}_{1}}\bigg[\mathbb{E}\iint_{Q_{0}}z^{2}dxdt+\mathbb{E}\iint_{Q}Z^{2}dxdt+\mathbb{E}\iint_{\Sigma}\widehat{Z}\,^{2}d\sigma dt\bigg],\end{aligned}$ (3.10)
where $\textbf{K}_{1}\equiv 1+\frac{1}{T}+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}$.
On the other hand, let $t\in(0,T)$, computing $d_{\tau}\|(z,z_{\Gamma})\|^{2}_{\mathbb{L}^{2}}$ by using the Itô’s formula. Then integrating the equality w.r.t $\tau\in(0,t)$ and taking the expectation on both sides, we conclude that
$\begin{aligned} &\,\mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma-\left(\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\right)\\ &=-2\mathbb{E}\int_{0}^{t}\int_{G}\mathcal{A}\nabla z\cdot\nabla zdxd\tau+2\mathbb{E}\int_{0}^{t}\int_{G}a_{1}z^{2}dxd\tau+2\mathbb{E}\int_{0}^{t}\int_{G}zB_{1}\cdot\nabla z\,dxd\tau\\ &\quad\,-\mathbb{E}\int_{0}^{t}\int_{G}Z^{2}dxd\tau-2\mathbb{E}\int_{0}^{t}\int_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma d\tau+2\mathbb{E}\int_{0}^{t}\int_{\Gamma}a_{2}z_{\Gamma}^{2}d\sigma d\tau\\ &\quad\,+2\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}B_{2}\cdot\nabla_{\Gamma}z_{\Gamma}\,d\sigma d\tau-\mathbb{E}\int_{0}^{t}\int_{\Gamma}\widehat{Z}^{2}d\sigma d\tau.\end{aligned}$
Recalling (1.1), and using Young’s inequality, it follows that for any $\varepsilon>0$,
$\begin{aligned} &\,\mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma-\left(\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\right)\\ &\leq-2\beta_{0}\mathbb{E}\int_{0}^{t}\int_{G}|\nabla z|^{2}dxd\tau+2\|a_{1}\|_{\infty}\mathbb{E}\int_{0}^{t}\int_{G}z^{2}dxd\tau+\varepsilon\mathbb{E}\int_{0}^{t}\int_{G}|\nabla z|^{2}\,dxd\tau\\ &\quad\,+\frac{\|B_{1}\|_{\infty}^{2}}{\varepsilon}\mathbb{E}\int_{0}^{t}\int_{G}z^{2}\,dxd\tau-\mathbb{E}\int_{0}^{t}\int_{G}Z^{2}dxd\tau-2\beta_{0}\mathbb{E}\int_{0}^{t}\int_{\Gamma}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma d\tau\\ &\quad\,+2\|a_{2}\|_{\infty}\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}^{2}d\sigma d\tau+\varepsilon\mathbb{E}\int_{0}^{t}\int_{\Gamma}|\nabla_{\Gamma}z_{\Gamma}|^{2}\,d\sigma d\tau+\frac{\|B_{2}\|_{\infty}^{2}}{\varepsilon}\mathbb{E}\int_{0}^{t}\int_{\Gamma}z^{2}_{\Gamma}\,d\sigma d\tau\\ &\quad\,-\mathbb{E}\int_{0}^{t}\int_{\Gamma}\widehat{Z}^{2}d\sigma d\tau.\end{aligned}$ (3.11)
Taking a small $\varepsilon$ in (3.11), we conclude that
$\begin{aligned} \mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma&\leq\,\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\\ &\hskip 11.38092pt+C\textbf{K}_{2}\left(\mathbb{E}\int_{0}^{t}\int_{G}z^{2}dxd\tau+\mathbb{E}\int_{0}^{t}\int_{\Gamma}z_{\Gamma}^{2}d\sigma d\tau\right),\end{aligned}$
where $\textbf{K}_{2}\equiv\|a_{1}\|_{\infty}+\|a_{2}\|_{\infty}+\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}$. Therefore, by Gronwall’s inequality, it follows that
$\begin{aligned} \mathbb{E}\int_{G}z^{2}(0)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(0)d\sigma\leq e^{CT\textbf{K}_{2}}\left(\mathbb{E}\int_{G}z^{2}(t)dx+\mathbb{E}\int_{\Gamma}z_{\Gamma}^{2}(t)d\sigma\right).\end{aligned}$ (3.12)
Finally, integrating (3.12) on $(T/4,3T/4)$ and combining the obtained inequality with (3.10), we conclude the observability inequality (3.8). ∎
4 Controllability Results: Proof of Theorems 1.2 and 1.3
This section is devoted to proving the null and approximate controllability results for the forward equation (1.7), as stated in Theorems 1.2 and 1.3, respectively. To proceed with the proofs, we first recall the following well-known result from functional analysis (see, e.g., [30] and [36, Theorem IV.2.2]).
Lemma 4.1.
Let $\mathbf{H}$, $\mathbf{U}$ and $\mathbf{Y}$ three Hilbert spaces and two bounded linear operators $M\in\mathcal{L}(\mathbf{H};\mathbf{Y})$ and $L\in\mathcal{L}(\mathbf{U};\mathbf{Y})$. Then we have
$\|M^{*}y\|_{\mathbf{H}}\leq\mathcal{C}\|L^{*}y\|_{\mathbf{U}}\,,\quad\forall y\in\mathbf{Y},\quad\textnormal{for some constant}\;\;\mathcal{C}>0,$ (4.1)
if and only if
$\mathcal{R}(M)\subset\mathcal{R}(L)\quad\textnormal{i.e.,}\quad\forall x_{0}\in\mathbf{H},\,\,\exists u\in\mathbf{U},\,\,\textnormal{s.t.}\,\,Mx_{0}=Lu,\,\quad\textnormal{and}\,\quad\|u\|_{\mathbf{U}}\leq\mathcal{C}\|x_{0}\|_{\mathbf{H}}.$ (4.2)
Moreover, the constant $\mathcal{C}$ in (4.1) and (4.2) is the same.
Let us first prove the approximate controllability of the equation (1.7).
Proof of Theorem 1.2.
Define the following controllability operator
$\mathcal{L}_{T}:L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))\longrightarrow L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2}),$
$\mathcal{L}_{T}(u,v_{1},v_{2})=(y(T,\cdot),y_{\Gamma}(T,\cdot)),$
where $(y,y_{\Gamma})$ is the solution of (1.7) with initial data $(y_{0},y_{\Gamma,0})=(0,0)$. It is easy to see that the equation (1.7) is approximately controllable at time $T$ if and only if $\mathcal{R}(\mathcal{L}_{T})$ is dense in $L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, equivalently the adjoint operator $\mathcal{L}_{T}^{*}$ is injective. Let us first compute $\mathcal{L}_{T}^{*}$, for this, multiplying the equations (3.1) and (1.7) using Itô’s formula, we find the following duality relation
$\begin{aligned} &\,\langle(y(T,\cdot),y_{\Gamma}(T,\cdot)),(z_{T},z_{\Gamma,T})\rangle_{L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})}-\langle(y_{0},y_{\Gamma,0}),(z(0,\cdot),z_{\Gamma}(0,\cdot))\rangle_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}\\ &=\mathbb{E}\iint_{Q}\mathbbm{1}_{G_{0}}uz\,dxdt+\mathbb{E}\iint_{Q}v_{1}Z\,dxdt+\mathbb{E}\iint_{\Sigma}v_{2}\widehat{Z}\,d\sigma dt.\end{aligned}$ (4.3)
It follows that for all $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$, the adjoint operator $\mathcal{L}_{T}^{*}$ is given by
[…]
Hence, (1.7) is approximately controllable at time $T$ if and only if the property
$(z,Z,\widehat{Z})=(0,0,0)\,\,\,\textnormal{in}\,\,\,Q_{0}\times Q\times\Sigma,\;\,\textnormal{a.s.}\;\;\Longrightarrow\;\,(z_{T},z_{\Gamma,T})=(0,0),\,\,\,\textnormal{in}\,\,\,G\times\Gamma,\;\,\textnormal{a.s.},$ (4.5)
holds. Since, the property (4.5) is achieved in Proposition 3.1, therefore the equation (1.7) is approximately controllable at time $T$. This completes the proof of Theorem 1.2. ∎
Let us now show the null controllability of (1.7).
Proof of Theorem 1.3.
Let us first define the following linear bounded operator
$\mathcal{M}_{T}:L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})\longrightarrow L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2}),\,\,\,\quad\mathcal{M}_{T}(y_{0},y_{\Gamma,0})=(y(T,\cdot),y_{\Gamma}(T,\cdot)),$
where $(y,y_{\Gamma})$ is the solution of (1.7) with controls $(u,v_{1},v_{2})\equiv(0,0,0)$. Also, we consider the following controllability operator
$\mathcal{L}_{T}:L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))\longrightarrow L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2}),$
$\mathcal{L}_{T}(u,v_{1},v_{2})=(y(T,\cdot),y_{\Gamma}(T,\cdot)),$
where $(y,y_{\Gamma})$ is the solution of (1.7) with initial data $(y_{0},y_{\Gamma,0})\equiv(0,0)$.
Recalling (4.3) for the duality relation between the equation (1.7) and its adjoint equation (3.1), it is easy to see that the adjoint operators of $\mathcal{M}_{T}$ and $\mathcal{L}_{T}$ are respectively defined by
[…]
for all $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$.
From the observability inequality (3.8), it is straightforward to see that for all $(z_{T},z_{\Gamma,T})\in L^{2}_{\mathcal{F}_{T}}(\Omega;\mathbb{L}^{2})$,
$\|\mathcal{M}_{T}^{*}(z_{T},z_{\Gamma,T})\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}\leq e^{CK}\|\mathcal{L}_{T}^{*}(z_{T},z_{\Gamma,T})\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))},$ (4.7)
where
$K\equiv 1+\frac{1}{T}+\|a_{1}\|_{\infty}^{2/3}+\|a_{2}\|_{\infty}^{2/3}+T(\|a_{1}\|_{\infty}+\|a_{2}\|_{\infty})+(1+T)(\|B_{1}\|_{\infty}^{2}+\|B_{2}\|_{\infty}^{2}).$
Let $(y_{0},y_{\Gamma,0})\in{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}$ be the initial state of (1.7). From Lemma 4.1 and the inequality (4.7), we deduce that $\mathcal{R}(\mathcal{M}_{T})\subset\mathcal{R}(\mathcal{L}_{T})$, and hence there exists controls $(\widehat{u},\widehat{v}_{1},\widehat{v}_{2})\in L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(G))\times L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))$ such that
$\|\widehat{u}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G_{0}))}+\|\widehat{v}_{1}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(G))}+\|\widehat{v}_{2}\|^{2}_{L^{2}_{\mathcal{F}}(0,T;L^{2}(\Gamma))}\leq e^{CK}\|(y_{0},y_{\Gamma,0})\|^{2}_{L^{2}_{\mathcal{F}_{0}}(\Omega;\mathbb{L}^{2})}.$ (4.8)
On the other hand, it is obvious that (1.7) is null controllable at time $T$ if and only if $\mathcal{R}(\mathcal{M}_{T})\subset\mathcal{R}(\mathcal{L}_{T})$. Therefore, the equation (1.7) is null controllable at time $T$, and from (4.8), the desired estimate (1.8) holds as well. This concludes the proof of Theorem 1.3. ∎
Appendix A Proof of Lemma 2.1
In this appendix, we follow some computations outlined in the proof of [6, Theorem 4.1] and [3, Lemma 3.2]. However, in our case, we must carefully handle the dependence of Carleman’s parameters on the final control time $T$, as well as work with the more general diffusion coefficients $\mathcal{A}$ and $\mathcal{A}_{\Gamma}$, instead of Laplacian operators.
First, we recall that the diffusion matrix operator $\mathcal{A}=(a^{jk})_{j,k=1,2,\cdots,N}$ satisfies the following assumptions:
  1. 1.
    $a^{jk}\in L^{\infty}_{\mathcal{F}}(\Omega;C^{1}([0,T];W^{2,\infty}(\overline{G})))$, and $a^{jk}=a^{kj}$ for all $j,k=1,2,\cdots,N$.
  2. 2.
    There exists a constant $\beta_{0}>0$ such that
    $\sum_{j,k=1}^{N}a^{jk}(\omega,t,x)\xi_{j}\xi_{k}\geq\beta_{0}|\xi|^{2},\quad\textnormal{for all}\;\;(\omega,t,x,\xi)\equiv(\omega,t,x,\xi_{1},\cdots,\xi_{N})\in\Omega\times(0,T)\times\overline{G}\times\mathbb{R}^{N}.$
Define a function $\ell\in C^{1,3}((0,T)\times\overline{G})$, and two auxiliary functions $\Psi\in C^{1,2}((0,T)\times\overline{G})$ and $\Phi\in C^{1,0}(\Sigma)$. For the sake of simplicity in notation, we let $y_{x_{j}}=\frac{\partial y}{\partial x_{j}}$ denote the partial derivative of the function $y$ with respect to the $j$-th component $x_{j}$ of a generic point $x=(x_{1},\cdots,x_{N})\in\mathbb{R}^{N}$. Similarly, we use $y_{x_{j}x_{k}}$ to denote the second partial derivative of $y$ with respect to the variables $x_{j}$ and $x_{k}$, and so on. Put
$\begin{cases}A=\displaystyle\sum_{j,k=1}^{N}(a^{jk}\ell_{x_{j}}\ell_{x_{k}}-a^{jk}_{x_{k}}\ell_{x_{j}}-a^{jk}\ell_{x_{j}x_{k}})-\Psi-\ell_{t},\\ B=2\Bigg[A\Psi+\displaystyle\sum_{j,k=1}^{N}(Aa^{jk}\ell_{x_{j}})_{x_{k}}\Bigg]-A_{t}+\sum_{j,k=1}^{N}(a^{jk}\Psi_{x_{k}})_{x_{j}},\\ c^{jk}=\displaystyle\sum_{j^{\prime},k^{\prime}=1}^{N}\Big[2a^{jk^{\prime}}(a^{j^{\prime}k}\ell_{x_{j^{\prime}}})_{x_{k^{\prime}}}-(a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}})_{x_{k^{\prime}}}\Big]+\frac{a^{jk}_{t}}{2}-\Psi a^{jk},\quad j,k=1,2,\cdots,N,\\ \widetilde{A}=\Phi-\ell_{t},\\ \widetilde{B}=-2\widetilde{A}\Phi-\widetilde{A}_{t}.\end{cases}$ (A.1)
From [27, Theorem 9.26], we have the following fundamental weighted identity for the stochastic parabolic-like operator $dz+\nabla\cdot(\mathcal{A}\nabla z)\,dt$.
Theorem A.1.
Let $z$ be an $H^{2}(G)$-valued Itô process. Set $\theta=e^{\ell}$ and $h=\theta z$. Then, for any $t\in[0,T]$ and a.e. $(x,\omega)\in\overline{G}\times\Omega$,
 $\begin{aligned} &\,2\theta\left[\nabla\cdot(\mathcal{A}\nabla h)+Ah\right]\left[dz+\nabla\cdot(\mathcal{A}\nabla z)dt\right]-2\nabla\cdot(\mathcal{A}\nabla h\,dh)\\ &+2\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dt\\ &=2\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dt+Bh^{2}dt+d\Bigg(-\sum_{j,k=1}^{N}a^{jk}h_{x_{j}}h_{x_{k}}+Ah^{2}\Bigg)\\ &\;\;\;\,+2\big[\nabla\cdot(\mathcal{A}\nabla h)+Ah\big]^{2}dt+\theta^{2}\sum_{j,k=1}^{N}a^{jk}(dz_{xj}+\ell_{x_{j}}dz)(dz_{x_{k}}+\ell_{x_{k}}dz)\\ &\;\;\;\,-\theta^{2}A(dz)^{2}.\end{aligned}$ (A.2)
We also have the following fundamental weighted identity for the stochastic parabolic-like operator $dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt$.
Theorem A.2.
Let $z_{\Gamma}$ be an $H^{2}(\Gamma)$-valued Itô process. Set $\theta=e^{\ell}$ and $h_{\Gamma}=\theta z_{\Gamma}$ with $z_{\Gamma}=z|_{\Gamma}$. Then, for any $t\in[0,T]$ and a.e. $(x,\omega)\in\Gamma\times\Omega$,
 $\begin{aligned} &\,2\theta\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt\right]-2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\,dh_{\Gamma})\\ &=\widetilde{B}h_{\Gamma}^{2}dt+d(-\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}+\widetilde{A}h_{\Gamma}^{2})+2[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}]^{2}dt\\ &\quad+\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}dt+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}-\theta^{2}\widetilde{A}|dz_{\Gamma}|^{2}-2\nabla_{\Gamma}\cdot(\Phi h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt\\ &\quad+2\Phi\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt+2h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi dt-2\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})dt-2\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zdt.\end{aligned}$ (A.3)
Proof.
We compute the first term on the left-hand side of (A.3). First, by the definition of $\theta$, we have
 $\begin{aligned} &\,2\theta\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[dz_{\Gamma}+\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})\,dt-\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+\widetilde{A}h_{\Gamma}dt+dh_{\Gamma}-\Phi h_{\Gamma}dt-\theta\partial_{\nu}^{\mathcal{A}}z\,dt\right]\\ &=2I(I_{1}+I_{2}+I_{3}+I_{4}),\end{aligned}$ (A.4)
where
$I=\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma},\quad I_{1}=\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+\widetilde{A}h_{\Gamma}dt,$
$I_{2}=dh_{\Gamma},\quad I_{3}=-\Phi h_{\Gamma}dt,\quad I_{4}=-\theta\partial_{\nu}^{\mathcal{A}}z\,dt.$
Firstly, we have
$2II_{1}=2\left[\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})+\widetilde{A}h_{\Gamma}\right]^{2}dt.$ (A.5)
By Itô’s formula, we derive that
$\begin{aligned} 2II_{2}=&2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}dh_{\Gamma})-d(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma})+(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt\\ &+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}+d(\widetilde{A}h_{\Gamma}^{2})-\widetilde{A}_{t}h_{\Gamma}^{2}dt-\widetilde{A}|dh_{\Gamma}|^{2},\end{aligned}$
which implies that
$\begin{aligned} 2II_{2}=&2\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}dh_{\Gamma})-d(\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma})+\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}dt\\ &+\mathcal{A}_{\Gamma}d\nabla_{\Gamma}h_{\Gamma}\cdot d\nabla_{\Gamma}h_{\Gamma}+d(\widetilde{A}h_{\Gamma}^{2})-\widetilde{A}_{t}h_{\Gamma}^{2}dt-\theta^{2}\widetilde{A}|dz_{\Gamma}|^{2}.\end{aligned}$ (A.6)
Similarly, we also obtain that
$2II_{3}=-2\nabla_{\Gamma}\cdot(\Phi h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma})dt+2\Phi\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}h_{\Gamma}dt+2h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi-2\widetilde{A}\Phi h_{\Gamma}^{2}dt.$ (A.7)
Finally, we have that
$2II_{4}=-2\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})dt-2\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zdt.$ (A.8)
Combining (A.4), (A.5), (A.6), (A.7), and (A.8), we deduce the desired identity (A.3). ∎
In what follows, we choose $\Psi=-2\sum_{j,k=1}^{N}a^{jk}\ell_{x_{j}x_{k}}$, $\Phi=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}(\mathcal{A}\nu\cdot\nu)|\nabla\ell|$ and $\ell$ is the function defined in (1.4). For a positive integer $n$, we denote by $O(\mu^{n})$ a function of order $\mu^{n}$ for large $\mu$, and similarly, we use the notation $O(e^{\mu|\psi|_{\infty}})$. By (1.4) and (1.5), it is easy to see that for any parameters $\lambda,\mu\geq 1$, we have for any $j,k=1,2,\cdots,N$,
$\ell_{t}=\lambda\alpha_{t},\quad\ell_{x_{j}}=\lambda\mu\varphi\psi_{x_{j}},\quad\ell_{x_{j}x_{k}}=\lambda\mu^{2}\varphi\psi_{x_{j}}\psi_{x_{k}}+\lambda\mu\varphi\psi_{x_{j}x_{k}},$ (A.9)
and also
$\begin{aligned} &\alpha_{t}=T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),\quad\varphi_{t}=T\varphi^{2}O(1),\\ &\alpha_{tt}=T^{2}\varphi^{3}O(e^{2\mu|\psi|_{\infty}}),\quad\varphi_{tt}=T^{2}\varphi^{3}O(1).\end{aligned}$ (A.10)
We have the following estimates for $A$, $B$, and $c^{jk}$ defined in (A.1) (see [3, Theorem A.2]).
Lemma A.1.
There exists a constant $C=C(G,G_{0},\mathcal{A},\mathcal{A}_{\Gamma})>0$ such that for sufficiently large $\mu\geq C$, and $\lambda\geq CT^{2}$, we have that
$A$ $=\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),$
$B\geq$  $\,2\beta_{0}^{2}\lambda^{3}\mu^{4}\varphi^{3}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu|\psi|_{\infty}}),$
$\sum_{j,k=1}^{N}c^{jk}\xi_{j}\xi_{k}\geq$ $\,\left[\beta_{0}^{2}\lambda\mu^{2}\varphi|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right]|\xi|^{2},\quad\forall\xi=(\xi_{1},\cdots,\xi_{N})\in\mathbb{R}^{N},$
for a.e. $(x,\omega)\in\overline{G\setminus G_{1}}\times\Omega$ and $t\in[0,T]$.
Similarly to the above lemma, for the quantities $\widetilde{A}$ and $\widetilde{B}$, we have the following result.
Lemma A.2.
The following identities hold:
$\widetilde{A}=-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}),$
$\widetilde{B}=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}),$
for all $x\in\Gamma\$, and for any $t\in[0,T]$,  a.s.
Proof.
We first have that
$\widetilde{A}$ $=\Phi-\ell_{t}$
$=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}(\mathcal{A}\nu\cdot\nu)|\nabla l|-\ell_{t}$
$=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi(\mathcal{A}\nu\cdot\nu)|\nabla\psi|+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
Notice that $\nabla\psi=\partial_{\nu}\psi\,\nu$ on $\Gamma$, then it follows that
$\widetilde{A}=\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi(\mathcal{A}\nu\cdot\nu)|\partial_{\nu}\psi|+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
Since $(\mathcal{A}\nu\cdot\nu)\partial_{\nu}\psi=\partial_{\nu}^{\mathcal{A}}\psi$ and that $\partial_{\nu}\psi<0$ on $\Gamma$, we conclude that
$\widetilde{A}=-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}}).$
On the other hand, we have that
$\widetilde{B}$ $=-2\widetilde{A}\Phi-\widetilde{A}_{t}$
$=-2\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi\right]$
$\quad\;+\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi_{t}\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi\varphi_{t}O(e^{2\mu|\psi|_{\infty}})$
$=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda\varphi^{2}O(\mu)+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}).$
Hence, we deduce that
$\widetilde{B}=-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}}).$
We are now in position to prove Lemma 2.1.
Proof of Lemma 2.1.
The proof is divided into three steps.
Step 1. Preliminary estimates.
Integrating (LABEL:5.121301) on $Q$, taking expectation on both sides and using the fact that $\theta(0,\cdot)=\theta(T,\cdot)=0$ and recalling (1.1), we conclude that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt-2\mathbb{E}\iint_{Q}\nabla\cdot(\mathcal{A}\nabla h\,dh)dx+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\\ &\quad\;+2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dxdt.\end{aligned}$ (A.11)
Integrating (A.3) on $\Sigma$, taking expectation on both sides and using the fact that $\theta(0,\cdot)=\theta(T,\cdot)=0$ and recalling (1.1), we conclude that
$\begin{aligned} &\,\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\Phi|\nabla_{\Gamma}h_{\Gamma}|^{2}d\sigma dt\\ &\leq\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi d\sigma dt\\ &\quad-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\\ &\quad+2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.\end{aligned}$ (A.12)
Combining (A.11) and (A.12), we get that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &+\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt+2\beta_{0}\mathbb{E}\iint_{\Sigma}\Phi|\nabla_{\Gamma}h_{\Gamma}|^{2}d\sigma dt\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\\ &\quad\;+2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\\ &\qquad\qquad\qquad\quad\;\;\;-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dxdt\\ &\quad\;-2\mathbb{E}\iint_{Q}\nabla\cdot(\mathcal{A}\nabla h\,dh)dx+\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi d\sigma dt\\ &\quad\;-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\\ &\quad\;+2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.\end{aligned}$
Hence, it follows that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt+\textbf{I}_{1}+\textbf{I}_{2}\\ &\leq\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+\sum_{i=3}^{9}\textbf{I}_{i}.\end{aligned}$ (A.13)
Step 2. Estimating the boundary terms $\textbf{I}_{i}$, $i=1,2,\cdots,9$.
From Lemma A.2, we first have that
$\begin{aligned} \textbf{I}_{1}=\mathbb{E}\iint_{\Sigma}\widetilde{B}h_{\Gamma}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\bigg[&-\frac{8}{\beta_{0}^{2}}\|\mathcal{A}\|^{2}_{\infty}\lambda^{2}\mu^{2}\varphi^{2}|\partial_{\nu}^{\mathcal{A}}\psi|^{2}+\lambda^{2}T\varphi^{3}O(\mu e^{2\mu|\psi|_{\infty}})\\ &+\lambda T^{2}\varphi^{3}O(e^{4\mu|\psi|_{\infty}})\bigg]z_{\Gamma}^{2}d\sigma dt,\end{aligned}$
which provides that
$\begin{aligned} \textbf{I}_{1}\geq&-C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt-C\lambda^{2}\mu Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &-C\lambda T^{2}e^{4\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.14)
We also have that
$\textbf{I}_{2}=2\beta_{0}\mathbb{E}\displaystyle\iint_{\Sigma}\Phi|\nabla_{\Gamma}v_{\Gamma}|^{2}d\sigma dt=4\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt,$
which leads to
$\textbf{I}_{2}=-4\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\displaystyle\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.15)
In what follows, we use the fact that for all $(t,x)\in\Sigma$, we have
$\nabla\ell=\lambda\mu\varphi\partial_{\nu}\psi\nu\quad\textnormal{and}\quad\nabla h=\theta\nabla z+\lambda\mu\theta\varphi z_{\Gamma}\partial_{\nu}\psi\nu.$
Notice that
$\textbf{I}_{3}$ $=2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)$
$\hskip 102.43008pt-\Psi a^{jk}h_{x_{j}}h+a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\Bigg]_{x_{k}}dxdt$
$=\textbf{I}_{3}^{1}+\textbf{I}_{3}^{2}+\textbf{I}_{3}^{3},$
where
$\textbf{I}_{3}^{1}=2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\Bigg[\sum_{j^{\prime},k^{\prime}=1}^{N}\big(2a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j^{\prime}}}h_{x_{j}}h_{x_{k^{\prime}}}-a^{jk}a^{j^{\prime}k^{\prime}}\ell_{x_{j}}h_{x_{j^{\prime}}}h_{x_{k^{\prime}}}\big)\Bigg]_{x_{k}}dxdt,$
$\textbf{I}_{3}^{2}=-2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}[\Psi a^{jk}h_{x_{j}}h]_{x_{k}}dxdt,$
$\textbf{I}_{3}^{3}=2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}\bigg[a^{jk}\Big(A\ell_{x_{j}}+\frac{\Psi_{x_{j}}}{2}\Big)h^{2}\bigg]_{x_{k}}dxdt.$
By direct computations, we obtain that
$\begin{aligned} \textbf{I}_{3}^{1}&=2\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+4\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\partial_{\nu}\psi\partial_{\nu}^{\mathcal{A}}\psi z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt\\ &\quad+4\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (A.16)
Using the inequity $4ab\leq\frac{8}{3}a^{2}+\frac{3}{2}b^{2}$ for the second term in the right-hand side of (A.16), we obtain that
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{5}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$
It follows that
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt.\end{aligned}$ (A.17)
To estimate the last two terms on the right-hand side of (A.17), we use the identity (see [1, Lemma 2.3])
$|\partial_{\nu}^{\mathcal{A}}z|^{2}-(\mathcal{A}\nabla_{\Gamma}z_{\Gamma}\cdot\nu)^{2}=(\mathcal{A}\nu\cdot\nu)(\mathcal{A}\nabla z\cdot\nabla z-\mathcal{A}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}).$
This implies that
$\begin{aligned} &2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt\\ &\leq-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi(\mathcal{A}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma})d\sigma dt,\end{aligned}$
which provides that
$\begin{aligned} &2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt-2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi(\mathcal{A}\nabla z\cdot\nabla z)(\mathcal{A}\nu\cdot\nu)d\sigma dt\\ &\leq-2\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.18)
From (A.17) and (A.18), we deduce that
$\begin{aligned} \textbf{I}_{3}^{1}&\leq-\frac{2}{3}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt+\frac{1}{2}\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad-2\lambda\mu\|\mathcal{A}\|_{\infty}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.19)
For large enough $\mu\geq C$, we have that
$\begin{aligned} \textbf{I}_{3}^{2}&=-2\mathbb{E}\iint_{\Sigma}\sum_{j,k=1}^{N}\Psi a^{jk}(\lambda\mu\theta\varphi z_{\Gamma}\partial_{\nu}\psi\nu_{j}+\theta z_{x_{j}})\theta z_{\Gamma}\nu_{k}d\sigma dt\\ &=-2\mathbb{E}\iint_{\Sigma}\Psi(\lambda\mu\theta^{2}\varphi z_{\Gamma}^{2}\partial_{\nu}\psi\mathcal{A}\nu\cdot\nu+\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}z)d\sigma dt\\ &=-2\mathbb{E}\iint_{\Sigma}\bigg(-2\lambda\mu^{2}\varphi\mathcal{A}\nabla\psi\cdot\nabla\psi-2\lambda\mu\varphi\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}x_{k}}\bigg)\\ &\hskip 56.9055pt(\lambda\mu\theta^{2}\varphi z_{\Gamma}^{2}\partial_{\nu}\psi\mathcal{A}\nu\cdot\nu+\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}z)d\sigma dt,\end{aligned}$
which leads to
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\lambda\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.20)
Applying Young’s inequality for the first term on the right-hand side of (A.20), we have that
$\begin{aligned} \textbf{I}_{3}^{2}\leq C\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+C\lambda^{2}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}d\sigma dt+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.21)
On the other hand, it is easy to see that
$\begin{aligned} \textbf{I}_{3}^{3}&=2\mathbb{E}\iint_{\Sigma}(A\lambda\mu\varphi\partial_{\nu}\psi\mathcal{A}\nu\cdot\nu+\frac{1}{2}\partial_{\nu}^{\mathcal{A}}\Psi)\theta^{2}z_{\Gamma}^{2}d\sigma dt\\ &=2\mathbb{E}\iint_{\Sigma}\left[\lambda^{3}\mu^{3}\theta^{2}\varphi^{3}z_{\Gamma}^{2}\partial_{\nu}\psi(\mathcal{A}\nu\cdot\nu)(\mathcal{A}\nabla\psi\cdot\nabla\psi)+\lambda^{2}\mu T\varphi^{3}O(e^{2\mu|\psi|_{\infty}})\right]d\sigma dt,\end{aligned}$
which implies for large enough $\lambda\geq CTe^{2\mu|\psi|_{\infty}}$,
$\begin{aligned} \textbf{I}_{3}^{3}\leq\frac{3}{2}\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}\partial_{\nu}\psi|\partial_{\nu}^{\mathcal{A}}\psi|^{2}z_{\Gamma}^{2}d\sigma dt.\end{aligned}$ (A.22)
We also have that
$\begin{aligned} \textbf{I}_{4}&=-2\mathbb{E}\iint_{Q}\nabla\cdot(\mathcal{A}\nabla h\,dh)dx\\ &=-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial_{\nu}\psi\theta^{2}z_{\Gamma}(\mathcal{A}\nu\cdot\nu)(-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})+\partial_{\nu}^{\mathcal{A}}z+F_{2}+\ell_{t}z_{\Gamma})d\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z(-\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})+\partial_{\nu}^{\mathcal{A}}z+F_{2}+\ell_{t}z_{\Gamma})d\sigma dt.\end{aligned}$
Then, it follows that
$\begin{aligned} \textbf{I}_{4}&=2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi\theta^{2}z_{\Gamma}\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}F_{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}^{2}\ell_{t}d\sigma dt\\ &\quad\;+2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt-2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}zF_{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}z\ell_{t}d\sigma dt,\end{aligned}$
which gives that
$\begin{aligned} \textbf{I}_{4}&=\mathbb{E}\iint_{\Sigma}(2\lambda\mu\theta^{2}\varphi\partial_{\nu}^{\mathcal{A}}\psi z_{\Gamma}+2\theta^{2}\partial_{\nu}^{\mathcal{A}}z)\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}F_{2}d\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\lambda\mu\varphi\partial^{\mathcal{A}}_{\nu}\psi\theta^{2}z_{\Gamma}^{2}\ell_{t}d\sigma dt-2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;-2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}zF_{2}d\sigma dt-2\mathbb{E}\iint_{\Sigma}\theta^{2}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}z\ell_{t}d\sigma dt.\end{aligned}$
Hence, it follows that
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}\psi||z_{\Gamma}||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||\nabla_{\Gamma}z_{\Gamma}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi||z_{\Gamma}||F_{2}|d\sigma dt\\ &\quad\;+2\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial^{\mathcal{A}}_{\nu}\psi|z_{\Gamma}^{2}|\ell_{t}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+2\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z||F_{2}|d\sigma dt+2\mathbb{E}\iint_{\Sigma}\theta^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z||\ell_{t}|d\sigma dt.\end{aligned}$ (A.23)
Notice that $|\ell_{t}|\leq C\lambda Te^{2\mu|\psi|_{\infty}}\varphi^{2}$, and using Young’s inequality in the right-hand side of (A.23), we find that
$\begin{aligned} \textbf{I}_{4}&\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}z_{\Gamma}^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt+C\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt+C\lambda^{2}\mu Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt\\ &\quad\;+CTe^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z_{\Gamma}|^{2}d\sigma dt.\end{aligned}$ (A.24)
We also have that
$\textbf{I}_{5}=\begin{aligned} \mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}\widehat{Z}^{2}d\sigma dt=\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]\widehat{Z}^{2}d\sigma dt,\end{aligned}$
which leads to
$\begin{aligned} \textbf{I}_{5}\leq C\lambda\mu T^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt+C\lambda Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.25)
Taking a large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, the inequality (A.25) implies hat
$\begin{aligned} \textbf{I}_{5}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.26)
By integration by parts, it is easy to see that
$\textbf{I}_{6}=-2\mathbb{E}\iint_{\Sigma}h_{\Gamma}\mathcal{A}_{\Gamma}\nabla_{\Gamma}h_{\Gamma}\cdot\nabla_{\Gamma}\Phi d\sigma dt$ $=\mathbb{E}\iint_{\Sigma}h_{\Gamma}^{2}\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}\Phi)d\sigma dt,$
which provides that
$\textbf{I}_{6}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi z_{\Gamma}^{2}dxdt.$ (A.27)
We also have that
$\textbf{I}_{7}=-\mathbb{E}\iint_{\Sigma}\theta^{2}(\mathcal{A}_{\Gamma})_{t}\nabla_{\Gamma}z_{\Gamma}\cdot\nabla_{\Gamma}z_{\Gamma}d\sigma dt\leq C\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.28)
It is not difficult to see that
$\textbf{I}_{8}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\partial_{\nu}^{\mathcal{A}}z\nabla_{\Gamma}\cdot(\mathcal{A}_{\Gamma}\nabla_{\Gamma}z_{\Gamma})d\sigma dt\leq\mathbb{E}\iint_{\Sigma}\theta^{2}|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt+\mathbb{E}\iint_{\Sigma}\theta^{2}|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt.$ (A.29)
On the other hand, we have that
$\textbf{I}_{9}=2\mathbb{E}\iint_{\Sigma}\theta^{2}\widetilde{A}z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt$ $=2\mathbb{E}\iint_{\Sigma}\theta^{2}\left[-\frac{2}{\beta_{0}}\|\mathcal{A}\|_{\infty}\lambda\mu\varphi\partial_{\nu}^{\mathcal{A}}\psi+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]z_{\Gamma}\partial_{\nu}^{\mathcal{A}}zd\sigma dt.$
It follows that
$\textbf{I}_{9}\leq C\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt+C\lambda Te^{2\mu|\psi|_{\infty}}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt,$
which implies that for large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$,
$\textbf{I}_{9}\leq C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}|z_{\Gamma}||\partial_{\nu}^{\mathcal{A}}z|d\sigma dt.$ (A.30)
Applying Young’s inequality for (A.30), we derive that
$\textbf{I}_{9}\leq C\lambda^{3}\mu^{2}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+C\lambda\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\partial_{\nu}^{\mathcal{A}}z|^{2}d\sigma dt.$ (A.31)
Combining (A.13), (A.15), (A.27), (A.28), (A.19), (A.21), (A.22), (A.24), (A.26), (A.14), (A.29), (A.31), and taking a large enough $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, we conclude that
$\begin{aligned} &2\mathbb{E}\iint_{Q}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}dxdt+\mathbb{E}\iint_{Q}Bh^{2}dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt+C\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt.\end{aligned}$ (A.32)
Step 3. Estimating the non-boundary terms.
From Lemma A.1, we have that for large $\mu\geq C$ and $\lambda\geq CT^{2}$,
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt=C\mathbb{E}\iint_{Q}\theta^{2}\left[\lambda^{2}\mu^{2}\varphi^{2}\sum_{j,k=1}^{N}a^{jk}\psi_{x_{j}}\psi_{x_{k}}+\lambda\varphi O(\mu^{2})+\lambda T\varphi^{2}O(e^{2\mu|\psi|_{\infty}})\right]Z^{2}dxdt.$
Taking a large $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$, it follows that
$C\mathbb{E}\iint_{Q}\theta^{2}AZ^{2}dxdt\leq C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt.$ (A.33)
Since $\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi(x)|^{2}>0$, by Lemma A.1, we obtain for a large $\mu\geq C$ and $\lambda\geq CT^{2}$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq 2\beta_{0}^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\Bigg[\left(\lambda\mu^{2}\varphi\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{2}+\lambda\varphi O(\mu)\right)|\nabla h|^{2}\\ &\quad+\left(\lambda^{3}\mu^{4}\varphi^{3}\min_{x\in\overline{G\setminus G_{1}}}|\nabla\psi|^{4}+\lambda^{3}\varphi^{3}O(\mu^{4})+\lambda^{2}T\varphi^{3}O(\mu^{2}e^{2\mu|\psi|_{\infty}})\right)h^{2}\Bigg]\,dxdt.\end{aligned}$ (A.34)
From (A.34), we conclude for a large $\mu\geq C$ and $\lambda\geq C(e^{2\mu|\psi|_{\infty}}T+T^{2})$,
 $\begin{aligned} &\,2\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\sum_{j,k=1}^{N}c^{jk}h_{x_{j}}h_{x_{k}}\,dxdt+\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}Bh^{2}\,dxdt\\ &\geq C\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G\setminus G_{1}}\varphi\left(|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\right)\,dxdt.\end{aligned}$ (A.35)
Notice that
$\frac{1}{C}\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)\leq|\nabla h|^{2}+\lambda^{2}\mu^{2}\varphi^{2}h^{2}\leq C\theta^{2}\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right).$ (A.36)
Combining (A.32), (A.33), (LABEL:5.232404) and (A.36), we conclude for large enough $\mu\geq C$ and $\lambda\geq C\left(e^{2\mu|\psi|_{\infty}}T+T^{2}\right)$,
$\begin{aligned} &\,\lambda\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &=\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\left(\int_{G\setminus G_{1}}+\int_{G_{1}}\right)\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt\\ &\quad\;+\lambda^{3}\mu^{3}\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{3}z_{\Gamma}^{2}d\sigma dt+\lambda\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi|\nabla_{\Gamma}z_{\Gamma}|^{2}d\sigma dt\\ &\leq C\mathbb{E}\iint_{Q}\theta^{2}F_{1}^{2}dxdt+C\mathbb{E}\iint_{\Sigma}\theta^{2}F_{2}^{2}d\sigma dt\\ &\quad\;+C\lambda^{2}\mu^{2}\mathbb{E}\iint_{Q}\theta^{2}\varphi^{2}Z^{2}dxdt+\lambda^{2}\mu\mathbb{E}\iint_{\Sigma}\theta^{2}\varphi^{2}\widehat{Z}^{2}d\sigma dt\\ &\quad\;+\lambda\mu^{2}\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi\left(|\nabla z|^{2}+\lambda^{2}\mu^{2}\varphi^{2}z^{2}\right)dxdt.\end{aligned}$ (A.37)
Finally, to eliminate the gradient term on the right-hand side of (A.37), we proceed exactly as in the proof of [27, Theorem 9.39] by choosing a cut-off function $\zeta\in C^{\infty}_{0}(G_{0};[0,1])$ with $\zeta\equiv 1$ in $G_{1}$, then computing $d(\theta^{2}\varphi\zeta^{2}z^{2})$ and integrating the equality over $Q$, and taking the expectation on both sides. We arrive at
$\mathbb{E}\int_{0}^{T}\int_{G_{1}}\theta^{2}\varphi|\nabla z|^{2}\,dxdt\leq C\mathbb{E}\iint_{Q_{0}}\theta^{2}\left(\frac{1}{\lambda^{2}\mu^{2}}F_{1}^{2}+\lambda^{2}\mu^{2}\varphi^{3}z^{2}\right)dxdt.$ (A.38)
Combining (A.37) and (A.38), we deduce the desired Carleman inequality (2.1). This concludes the proof of Lemma 2.1. ∎